### v3

# Directional testing for one-way MANOVA in divergent dimensions

Abstract
Testing the equality of mean vectors across $g$ different groups plays an important role in many scientific fields. In regular frameworks, likelihood-based statistics under the normality assumption offer a general solution to this task. However, the accuracy of standard asymptotic results is not reliable when the dimension $p$ of the data is large relative to the sample size $n_{i}$ of each group. We propose here an exact directional test for the equality of $g$ normal mean vectors with identical unknown covariance matrix in a high dimensional setting, provided that $\sum_{i=1}^{g}n_{i}\geq p+g+1$. In the case of two groups ($g=2$), the directional test coincides with the Hotelling’s $T^{2}$ test. In the more general situation where the $g$ independent groups may have different unknown covariance matrices, although exactness does not hold, simulation studies show that the directional test is more accurate than most commonly used likelihood-based solutions, at least in a moderate dimensional setting in which $p=O(n_{i}^{\tau})$, $\tau\in(0,1)$. Robustness of the directional approach and its competitors under deviation from the assumption of multivariate normality is also numerically investigated. Our proposal is here applied to data on blood characteristics of male athletes and to microarray data storing gene expressions in patients with breast tumors.
Abstract
Section S1 provides auxiliary computational results for the correction factor $\gamma$ in skovgaard:2001’s modifications, as well as the observed information matrix operator used for testing hypothesis (8) of the main text. Section S2 presents the proofs for Lemma 1, Theorem 1 and Proposition 1. Section S3 reports the extensive simulation studies for the homoscedastic one-way MANOVA; Specifically, Sections S3.1–S3.4 examine varing dimenions $p$ and the group numbers $g$. While the directional test maintains robust performance, the log-likelihood ratio test fails even at small $p$. Other chi-square approximations, i.e. Bartlett correction and two skovgaard:2001’s modifications, deteriorate as $p$ increases, particularly in the extreme settings of Section S3.4. Sections S3.5–S3.6 assess robustness under model misspecification for high and moderate dimensional setups. Furthermore, Section S4 reports extensive simulations for the heteroscedastic one-way MANOVA. Finally, Section S5 provides results from an additional application to breast cancer microarray data.
Keywords: Behrens-Fisher problem, High dimension, Likelihood ratio test, Model misspecification, Multivariate normal distribution.
1 Introduction
Hypothesis testing for multivariate mean vectors is a very important inferential problem in many applied research fields. Likelihood-based statistics and usual asymptotic results offer a general solution to this task in parametric models. Typically, such solutions are accurate when the model dimension $p$ and the sample size $n$ match the standard asymptotic setting, where the dimension of the parameter is considered fixed as the sample size increases. However, usual asymptotic results are no longer guaranteed when $p$ is not negligible with respect to $n$ (see for instance jiang:2013; sur2019likelihood; tang2020modified; He2020B). As a simple illustration, the left panel of Figure 1 shows the empirical null distribution based on 10,000 Monte Carlo samples of $p$-values obtained using the asymptotic chi-square approximation for the likelihood ratio statistic when testing the equality of $g=4$ normal mean vectors under the assumption of common unknown covariance matrix. This problem is known as homoscedastic one-way multivariate analysis of variance (MANOVA), and the simulation setup is taken from the Pottery data in the R (Rcode) package car (car): the group sizes are $5,2,5$ and $14$, respectively, with $n=\sum_{i=1}^{4}n_{i}=26$, and the dimension of the vectors is $p=5$. It is clear from Figure 1 that the standard approximate $p$-values obtained from the likelihood ratio statistic are far from being uniform, as opposed to the directional $p$-values proposed in this paper and shown in the right panel.
Figure 1: Empirical null distribution of $p$-values from the likelihood ratio test (left) and directional test (right) for the hypothesis of equality of normal mean vectors in $g=4$ groups with identical covariance matrix, based on $10,000$ Monte Carlo simulations. Data are generated from a $N_{5}(\mu,\Lambda^{-1})$ distribution with mean vector $\mu$ and covariance matrix $\Lambda^{-1}$ equal to the sample mean and sample covariance matrix, respectively, of the Pottery dataset. The total sample size is $n=\sum_{i=1}^{4}n_{i}=26$.
Settings like the former, in which the values of $p$ and $n$ are comparable, may be framed in a $(p,n)$-divergent dimensional asymptotic setting where both $p$ and $n$ are allowed to increase. Indeed, the data dimension $p$ is related to the number of parameters, e.g., for the homoscedastic one-way MANOVA case in Section 3.1, the numbers of parameters is $gp+p(p+1)/2$. Inspired by battey2022some, we distinguish between three asymptotic regimes: moderate dimensional, high dimensional and ultra-high dimensional. Here we do not deal with the ultra-high dimensional asymptotic regime, in which $p/n$ diverges or tends to a limit greater than one. Instead, we focus on the moderate dimensional asymptotic regime, in which $p/n\to 0$, in particular with $p=O(n^{\tau})$, $\tau\in(0,1)$, and on the high dimensional asymptotic regime, in which $p/n\to\kappa\in(0,1)$. In the moderate dimensional asymptotic regime, He2020B proved that in testing the equality of normal mean vectors with identical covariance matrix the likelihood ratio test is valid if $p/n^{2/3}\to 0$, while the analogous condition for its Bartlett corrected test is $p/n^{4/5}\to 0$. To our knowledge, no similar results are available for the heteroscedastic case.
Higher-order likelihood solutions based on saddlepoint approximations might generally give substantial improvements over first-order solution, especially in high dimensions (see, e.g., tang2020modified). Among these, directional inference on a vector parameter of interest, as developed by sartori:2014 and sartori:2016, has proven to be particularly accurate when testing canonical parameters in exponential families. Its accuracy descends from that of the underlying saddlepoint approximation to the conditional density of the canonical statistic of interest. Empirical results in sartori:2014 and sartori:2016 showed that directional tests are extremely accurate even in settings where the dimension of the parameter of interest, although lower than the sample size, has a comparable order. The use of a saddlepoint approximation, indeed, guarantees a fairly constant relative error.
The use of a directional test may be motivated by the fact that, in standard asymptotics, the directional test is first-order equivalent to the likelihood ratio test. Yet, if first-order approximations are needed for the distribution of the latter, the directional test may be more convenient, given its improved accuracy (skogaard:1988; sartori2017introduction). Moreover, sartori:2019Ftest showed that the directional test coincides with many well-known exact tests. For example, when testing a specific value for the mean of a multivariate normal distribution, the directional test coincides with the exact Hotelling’s $T^{2}$ test. huang2021directional found other instances in which the directional $p$-value is exactly uniformly distributed. Such examples are related to several prominent inferential problems in which the independence of components or the equality of covariance matrices is tested in high dimensional multivariate normal models. Finally, claudia2023ss showed the accurate properties of directional tests for covariance selection in high dimensional Gaussian graphical models.
Concentrating on tests for the hypothesis of equality of mean vectors in $g$ independent groups, this work makes a number of contributions to the current literature. First, under the assumptions of normality and identical unknown covariance matrix, we prove that the directional $p$-value is exactly uniformly distributed provided that $n=\sum_{i=1}^{g}n_{i}\geq p+g+1$, and coincides with the Hotelling’s $T^{2}$ test when $g=2$. For the more general case with $g$ unknown group covariance matrices, known as the Behrens-Fisher problem if $g=2$, the directional test is not exact. Still, we show by means of extensive simulation studies that the directional approach overperforms standard first-order solutions as well as other higher-order modifications (skovgaard:2001) in moderate dimensional settings. In addition, the robustness of the available solutions to the normality assumption is empirically investigated, considering multivariate $t$, skew-normal or Laplace true generating processes in the Supplementary Material. Finally, the proposed directional test is applied for the analysis of two datasets: one from a study on the blood characteristics of male athletes, and another, presented in the Supplementary Material, from a microarray analysis on breast cancer patients.
The rest of the paper is organized as follows. Section 2 presents some background information. In particular, Section 2.1 reviews some important likelihood-based statistics in exponential families, Section 2.2 reports the steps to compute the directional $p$-value, and Section 2.3 details the necessary quantities for the multivariate normal model. The main results in Section 3 are for hypotheses concerning: (i) the equality of $g$ normal mean vectors with identical covariance matrix (Section 3.1); (ii) the equality of $g$ normal mean vectors with different covariance matrices (Section 3.2). For hypothesis (i), we prove the exact uniform distribution of the directional $p$-value under the null. In Section 4 we report empirical results under the assumed model. An application of the considered methodology to real data, specifically blood characteristics of male athletes, is illustrated in Section 5, while Section 6 concludes with a discussion. The Supplementary Material contains auxiliary computational results, proofs of the theorems, additional simulation studies, including larger group sizes and robustness analyses, as well as the second application involving breast cancer patient microarray data.
2 Background
2.1 Notation and setup
Suppose that data $y=(y_{1},\dots,y_{n})^{\top}$ are generated from the model $f(y;\theta)$ with parameter $\theta=(\theta_{1},\dots,\theta_{q})^{\top}$. The log-likelihood function is $\ell(\theta)=\ell(\theta;y)=\log f(y;\theta)$, and we are interested in testing the null hypothesis $H_{\psi}:\psi(\theta)=\psi$ on the $d$-dimensional parameter of interest $\psi$. It will often be the case that $\psi$ is a component of $\theta$, possibly after a reparameterization. Assume the partition $\theta=(\psi^{\top},\lambda^{\top})^{\top}$ holds, with $\lambda$ a (qd)(q-d)-dimensional nuisance parameter. Let $\hat{\theta}$ denote the maximum likelihood estimate of $\theta$ and $\hat{\theta}_{\psi}$ its constrained maximum likelihood estimate under $H_{\psi}$, i.e., $\hat{\theta}_{\psi}=(\psi^{\top},\hat{\lambda}_{\psi}^{\top})^{\top}$.
Several likelihood-based statistics can be used to test the hypothesis $H_{\psi}$. The likelihood ratio statistic, a parameterization-invariant measure, is
$W=2\{\ell(\hat{\theta})-\ell(\hat{\theta}_{\psi})\}.$ (1)
When $q$ is fixed and $n\to\infty$, the statistic $W$ has a $\chi^{2}_{d}$ asymptotic null distribution with relative error of order $O(n^{-1})$. A correction of $W$ proposed by bartlett:1937 rescales the likelihood ratio statistic by its expectation under $H_{\psi}$, that is
$W_{BC}=\frac{d}{E(W)}W,$ (2)
and follows asymptotically a $\chi^{2}_{d}$ null distribution with relative error of order $O(n^{-2})$ (mccullage:2018, Section 7.4). More details on the expectation $E(W)$ can be found in huang2021directional.
skovgaard:2001 introduced two improvements on $W$, namely
$W^{*}=W\left(1-\frac{\log\gamma}{W}\right)^{2}\quad\text{and}\quad W^{**}=W-2\log\gamma,$ (3)
which also have approximate $\chi^{2}_{d}$ distributions when the null hypothesis holds. The correction factor $\gamma$ can be found in skovgaard:2001, and is also reported in S1.1 with the notation used here.
The test statistics presented so far are omnibus measures of departure of the data from $H_{\psi}$: their $p$-value results from averaging the deviations from $H_{\psi}$ in all the potential directions of the parameter space. We will now introduce the directional test developed by sartori:2014 and sartori:2016, which measures the departure from $H_{\psi}$ along the direction indicated by the observed data.
2.2 Directional testing
Let $\varphi=\varphi(\theta)$ be a reparameterization of the original model. Suppose we have an exponential family model with sufficient statistic $u=u(y)$ and canonical parameter $\varphi$, with density
$f(y;\theta)=\exp\left[\varphi(\theta)^{\top}u-K\{\varphi(\theta)\}\right]h(y),$ (4)
[…]
$\ell(\varphi;s)=\varphi(\theta)^{\top}s+\ell^{0}(\theta),$
where $\ell^{0}(\theta)=\ell(\theta;y^{0})$ is the observed log-likelihood function. The saddlepoint approximation (see, e.g., salvan:1997, Section 10) to the exponential model on $\mathbb{R}^{q}$ is
$f(s;\varphi)=c\exp\{\ell(\varphi;s)-\ell(\hat{\varphi};s)\}|J_{\varphi\varphi}(\hat{\varphi})|^{-1/2},$
[…]
The hypothesis $H_{\psi}$ specifies a value for the parameter $\psi=\psi(\varphi)$. Following sartori:2016, the reduced model in $\mathbb{R}^{d}$ is given by
$h(s;\psi)=c\exp\{\ell(\hat{\varphi}_{\psi};s)-\ell(\hat{\varphi};s)\}|J_{\varphi\varphi}(\hat{\varphi})|^{-1/2}|\tilde{J}_{(\lambda\lambda)}|^{1/2},\quad s\in L_{\psi}^{0},$ (5)
where $L_{\psi}^{0}$ is the $d$-dimensional plane obtained by setting $\hat{\lambda}_{\psi}=\hat{\lambda}_{\psi}^{0}$, and the observed information for the nuisance parameter has been recalibrated to $\varphi$ as follows:
$|\tilde{J}_{(\lambda\lambda)}|=|J_{\lambda\lambda}(\hat{\varphi}_{\psi};s)||\partial\varphi(\theta)/\partial\lambda|^{-2}_{\hat{\theta}_{\psi}}.$ (6)
The directional test is constructed by defining a line $L_{\psi}^{*}$ through the observed value $s^{0}=0_{q}$ and the expected value of $s$ under $H_{\psi}$ which depends on the observed data point $y^{0}$, i.e. $s_{\psi}=-\ell_{\varphi}^{0}\left(\hat{\varphi}^{0}_{\psi}\right)=-\partial\ell^{0}\left\{\theta(\varphi)\right\}/\partial\varphi\left|{}_{\varphi=\hat{\varphi}^{0}_{\psi}}\right$.. We parameterize this line by $t\in\mathbb{R}$, namely $s(t)=s_{\psi}+t(s^{0}-s_{\psi})$. In particular, $s(0)=s_{\psi}$, corresponding to $H_{\psi}$, and $s(1)=s^{0}$, corresponding to the observed data. Then, the directional test for $s$ is conditional on being on the line $L^{*}_{\psi}$, i.e. conditional on s(t)/s(t)s(t)/$|$s(t)$|$. The directional $p$-value measuring the departure from $H_{\psi}$ along the line $L_{\psi}^{*}$ is defined as the probability that $s(t)$ is as far or farther from $s_{\psi}$ than is the observed value $s^{0}$:
$p(\psi)=\frac{\int_{1}^{t_{\text{sup}}}{t^{d-1}h\{s(t);\psi\}}\text{d}t}{\int_{0}^{t_{\text{sup}}}{t^{d-1}h\{s(t);\psi\}}\text{d}t},$ (7)
where the denominator is a normalizing constant. The upper limit $t_{\text{sup}}$ of the integrals in (7) is the largest value of $t$ for which the maximum likelihood estimate $\hat{\varphi}(t)$, corresponding to $s(t)$, exists. See sartori:2016 for more details.
In the particular case where $\psi$ and $\lambda$ are linear functions of the canonical parameter of an exponential family, the quantity (6) does not depend on $s$ and can therefore be ignored in computing the directional $p$-value (7). Moreover, the expected value $s_{\psi}$ simplifies to $s_{\psi}=\left\{-\ell_{\psi}^{0}\left(\hat{\varphi}^{0}_{\psi}\right)^{\top},{{0}_{q-d}^{\top}}\right\}^{\top}$.
If the original model $f(y;\theta)$ has not the exponential family form, then a tangent exponential model is used instead of (4). The construction of the tangent exponential model and its saddlepoint approximation are described in sartori:2016; see also davison2022tangent. Here this step is not needed since we work with the normal model which belongs to the exponential family, yet the saddlepoint approximation (5) will be used when $\psi$ is not a linear function of the canonical parameter, as in the heteroscedastic one-way MANOVA of Section 3.2.
2.3 Independent groups from multivariate normal distributions
Consider $g$ independent groups, and denote by $y_{ij}$ the independent observations from the $i$th group with multivariate normal distribution $N_{p}(\mu_{i},\Lambda_{i}^{-1})$, $i\in\{1,\dots,g\};j\in\{1,\dots,n_{i}\}$. The mean vectors $\mu_{i}$ and the concentration matrices $\Lambda_{i}$, symmetric and positive definite, are assumed unknown. Let $\text{tr}(A)$ indicate the trace of a square matrix $A$, and $\text{vec}(A)$ be the vector stacking the columns of $A$ one by one. We also define the vector $\text{vech}(A)$, obtained from $\text{vec}(A)$ by eliminating all upper triangular elements of $A$ when this is symmetric. These two vectors satisfy the relationship $D_{p}\text{vech}(A)=\text{vec}(A)$, where $D_{p}$ is the so-called duplication matrix (magnus1999, Section 3.8).
We rewrite the data from the $i$th group as $y_{i}=[y_{i1}\cdots y_{in_{i}}]^{\top}$, which is a $n_{i}\times p$ matrix. Then, the log-likelihood for the parameter $\theta=\{\mu_{1}^{\top},\dots,\mu_{g}^{\top},\text{vech}(\Lambda_{1}^{-1})^{\top},\dots,\text{vech}({\Lambda}_{g}^{-1})^{\top}\}^{\top}$ is
$\ell(\theta)$ $=\sum_{i=1}^{g}n_{i}\mu_{i}^{\top}\Lambda_{i}\bar{y}_{i}-\frac{1}{2}{\text{tr}}(\Lambda_{i}y_{i}^{\top}y_{i})+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\mu_{i}^{\top}\Lambda_{i}\mu_{i},$
where $\bar{y}_{i}=y_{i}^{\top}1_{n_{i}}/n_{i}$ with $1_{n_{i}}$ a $n_{i}$-dimensional vector of ones, $i\in\{1,\dots,g\}$. In this exponential family model, the canonical parameter is $\varphi=\{\xi_{1}^{\top},\dots,\xi_{g}^{\top},$ ${\text{vech}}(\Lambda_{1})^{\top},\dots,{\text{vech}}(\Lambda_{g})^{\top}\}^{\top}=\{\mu_{1}^{\top}\Lambda_{1},\dots,\mu_{g}^{\top}\Lambda_{g},$ ${\text{vech}}(\Lambda_{1})^{\top},\dots,{\text{vech}}(\Lambda_{g})^{\top}\}^{\top}$. The log-likelihood as a function of $\varphi$ is $\ell(\varphi)=\sum_{i=1}^{g}\ell_{i}(\varphi_{i})$ with $\varphi_{i}=\{\xi_{i}^{\top},{\text{vech}}(\Lambda_{i})^{\top}\}^{\top}$, and the $i$-th group’s contribution is
$\ell_{i}(\varphi_{i})$ $=n_{i}\xi_{i}^{\top}\bar{y}_{i}-\frac{1}{2}{\text{tr}}(\Lambda_{i}y_{i}^{\top}y_{i})+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\xi_{i}^{\top}\Lambda_{i}^{-1}\xi_{i}$
$=\xi_{i}^{\top}n_{i}\bar{y}_{i}-\text{vech}(\Lambda_{i})^{\top}\left\{\frac{1}{2}D_{p}^{\top}D_{p}\text{vech}(y_{i}^{\top}y_{i})\right\}+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\xi_{i}^{\top}\Lambda_{i}^{-1}\xi_{i}.$
The score function is $\ell_{\varphi}(\varphi)=\left.{\partial}\ell(\varphi)\right/{\partial\varphi}=\{\ell_{\varphi_{1}}(\varphi_{1}),\dots,\ell_{\varphi_{g}}(\varphi_{g})\}^{\top}$, with
$\ell_{\varphi_{i}}(\varphi_{i})$ $=\left\{\ell_{\xi_{i}}(\varphi_{i})^{\top},\ell_{\text{vech}(\Lambda_{i})}(\varphi_{i})^{\top}\right\}^{\top}$
$=\left\{n_{i}\bar{y}_{i}^{\top}-n_{i}\xi_{i}^{\top}\Lambda_{i},\;\frac{n_{i}}{2}\text{vech}(\Lambda_{i}^{-1}-y_{i}^{\top}y_{i}/n_{i}+\Lambda_{i}^{-1}\xi_{i}\xi_{i}^{\top}\Lambda_{i}^{-1})^{\top}\right\}^{\top}.$
The maximum likelihood estimates for $\mu_{i}$ and $\Lambda_{i}^{-1}$ are $\hat{\mu}_{i}=\bar{y}_{i}$ and $\hat{\Lambda}_{i}^{-1}=y_{i}^{\top}y_{i}/n_{i}-\bar{y}_{i}\bar{y}_{i}^{\top}$, respectively; hence, $\hat{\xi}_{i}=\hat{\Lambda}_{i}\hat{\mu}_{i}$, $i\in\{1,\dots,g\}$. Moreover, the observed information matrix $J_{\varphi\varphi}(\varphi)$ can be written in a block-diagonal form, with the group-specific diagonal block $J_{\varphi_{i}\varphi_{i}}(\varphi_{i})$, $i\in\{1,\dots,g\}$, equal to (huang2021directional, Section 2.3)
$J_{\varphi_{i}\varphi_{i}}(\varphi_{i})=\left[\begin{array}[]{cc}n_{i}\Lambda_{i}^{-1}&\quad-n_{i}(\xi_{i}^{T}\Lambda_{i}^{-1}\otimes\Lambda_{i}^{-1})D_{p}\\ -n_{i}D_{p}^{T}(\Lambda_{i}^{-1}\xi_{i}\otimes\Lambda_{i}^{-1})&\quad\frac{n}{2}D_{p}^{T}\{\Lambda_{i}^{-1}(I_{p}+2\xi_{i}\xi_{i}^{T}\Lambda_{i}^{-1})\otimes\Lambda_{i}^{-1}\}D_{p}\end{array}\right],$ (10)
where $\otimes$ denotes the Kronecker product (see, e.g., Lauritzen:1996, Section 5.1). Then, the determinant of $J_{\varphi\varphi}$ is such that $|J_{\varphi\varphi}(\varphi)|=\prod_{i=1}^{g}|J_{\varphi_{i}\varphi_{i}}(\varphi_{i})|\propto\prod_{i=1}^{g}|{\Lambda}_{i}^{-1}|^{p+2}$.
If the covariance matrices of the $g$ groups are the same, $\Lambda^{-1}_{i}=\Lambda^{-1}$ for $i\in\{1,\dots,g\}$, the canonical parameter is $\varphi=\{\xi_{1},\dots,\xi_{g},$ $\text{vech}(\Lambda)^{\top}\}^{\top}=\left\{\mu_{1}^{\top}\Lambda,\dots,\mu_{g}^{\top}\Lambda,\text{vech}(\Lambda)^{\top}\right\}^{\top}$, and the maximum likelihood estimate for $\mu_{i}$ and $\Lambda^{-1}$ are, respectively, $\hat{\mu}_{i}=\bar{y}_{i}$ and $\hat{\Lambda}^{-1}=B/n$ with $B=\sum_{i=1}^{g}y_{i}^{\top}y_{i}-n_{i}\bar{y}_{i}\bar{y}_{i}^{\top}$ and $n=\sum_{i=1}^{g}n_{i}$. In this setting, the observed information matrix $J_{\varphi\varphi}(\varphi)$ can be computed in block form (see Section S1.2).
3 One-way MANOVA problems
3.1 Homoscedastic one-way MANOVA
Suppose that $y_{ij}$ are independent observations from distributions $N_{p}(\mu_{i},\Lambda^{-1})$, for $i\in\{1,\dots,g\}\;(g\geq 2)$ and $j\in\{1,\dots,n_{i}\}$. We are interested in testing the equality of the $g$ mean vectors:
$H_{\psi}:\mu_{1}=\dots=\mu_{g}.$ (11)
The hypothesis problem (11) is equivalent to testing
$H_{\psi}:\Lambda\mu_{1}=\dots=\Lambda\mu_{g}.$
In this framework, the $q$-dimensional canonical parameter is $\varphi=\{\xi_{1}^{\top},\dots,\xi_{g}^{\top},{\text{vech}}(\Lambda^{\top})\}^{\top}=\{\mu_{1}^{\top}\Lambda,\dots,\mu_{g}^{\top}\Lambda,$ ${\text{vech}}(\Lambda^{\top})\}^{\top}$, with $d$-dimensional parameter of interest $\psi=(\xi_{2}^{\top}-\xi_{1}^{\top},\dots,\xi_{g}^{\top}-\xi_{1}^{\top})^{\top}$ and (qd)(q-d)-dimensional nuisance parameter $\lambda=\{\xi_{1}^{\top},$ ${\text{vech}}(\Lambda^{\top})\}^{\top}$. The parameter of interest $\psi$ is therefore a linear function of the canonical parameter $\varphi$. The maximum likelihood estimates of parameters $\mu_{i}$ and $\Lambda^{-1}$ are given in Section 2.3. The constrained maximum likelihood estimate are $\hat{\mu}_{0}=\bar{y}$ and $\hat{\Lambda}^{-1}_{0}=\left(\sum_{i=1}^{g}y_{i}^{\top}y_{i}-n\bar{y}\bar{y}^{\top}\right)/n=(A+B)/n$ with $\bar{y}=\sum_{i=1}^{g}n_{i}\bar{y}_{i}/n$, $A=\sum_{i=1}^{g}n_{i}\bar{y}_{i}^{\top}\bar{y}_{i}-n_{i}\bar{y}\bar{y}^{\top}$, and $B$ as defined in Section 2.3.
There are several likelihood-based tests for the hypothesis problem (11) when the dimension $p$ is fixed and the group size $n_{i}$ goes to infinity. However, as discussed in Section 1, our focus here is on improving the existing standard approximations when $p$ is relatively large with respect to $n_{i}$. Below we provide the necessary and sufficient conditions for the validity of the directional test in the high dimensional regime where $p/n_{i}\to\kappa\in(0,1),i\in\{1,\dots,g\}$.
First, we summarize here the key methodological steps to compute the directional $p$-value (7) for testing hypothesis (11) (see also sartori:2014, for more details). Under $H_{\psi}$, we define the expected value of $s$ under $H_{\psi}$ as $s_{\psi}=(n_{1}\bar{y}^{\top}-n_{1}\bar{y}_{1}^{\top},\dots,$ $n_{g}\bar{y}^{\top}-n_{g}\bar{y}_{g}^{\top},0_{q-d})^{\top}$, and the line $s(t)=(1-t)s_{\psi}$. The tilted log-likelihood along the line $s(t)$ is then
$\ell(\varphi;t)=\sum_{i=1}^{g}\ell_{i}(\varphi_{i};t)=\sum_{i=1}^{g}\ell_{i}(\varphi_{i})+\varphi^{\top}_{i}s_{i}(t),$
where
$\ell_{i}(\varphi_{i};t)=$ ${n_{i}}\xi_{i}^{\top}\left\{(1-t)\bar{y}+t\bar{y}_{i}\right\}-\frac{1}{2}{\text{tr}}\left(\Lambda y_{i}^{\top}y_{i}\right)+\frac{n_{i}}{2}\log|\Lambda|-\frac{n_{i}}{2}\xi_{i}^{\top}\Lambda^{-1}\xi_{i}.$
The corresponding saddlepoint approximation is
$h\{s(t);\psi\}=$ $c\exp\left\{\frac{(n-p-g-1)}{2}\log|\hat{\Lambda}(t)|\right\},$
where $c$ is a normalizing constant. The following lemma states that $t_{\text{sup}}$ in (7), i.e. the largest $t$ such that $\hat{\Lambda}(t)^{-1}$ is positive definite, is equal to $1/\sqrt{{\nu_{(p)}}}$, where $\nu_{(p)}$ is the largest eigenvalue of the matrix $(B_{0}^{\top})^{-1}(A/n)B_{0}^{-1}$ with $B_{0}$ the square root of $\hat{\Lambda}_{0}^{-1}$, namely $\hat{\Lambda}^{-1}_{0}=B_{0}^{\top}B_{0}$. The function chol() in the R package Matrix (Matrix) can be used to compute $B_{0}$.
Lemma 1.
The estimator $\hat{\Lambda}^{-1}(t)$ is positive definite if and only if $t\in\left[0,1/\sqrt{\nu_{(p)}}\right]$, where $\nu_{(p)}$ is the largest eigenvalue of $(B_{0}^{\top})^{-1}(A/n)B_{0}^{-1}$, with $\hat{\Lambda}^{-1}_{0}=B_{0}^{\top}B_{0}$.
Thanks to the favorable properties of the multivariate normal distribution, we can show that the saddlepoint approximation to the conditional density of $s$ along the line $s(t)$ is exact, and so when the null hypothesis holds the directional $p$-value is exactly uniformly distributed even in the high dimensional asymptotic regime. The condition for the validity of this result is given in the following theorem.
 Theorem 1.
Assume that $p=p_{n}$ is such that $n\geq p+g+1$, with $n=\sum_{i=1}^{g}n_{i}$ and fixed $g$. Then, under the null hypothesis (11), the directional $p$-value is exactly uniformly distributed.
When the number of independent groups is $g=2$, the Hotelling’s $T^{2}$ test statistic with exact $F$ distribution can be used for hypothesis (11). We also prove that in this case the directional $p$-value coincides with the one from the Hotelling’s $T^{2}$ test.
 Proposition 1.
When $g=2$, the directional test is equivalent to the Hotelling’s $T^{2}$ test.
The proofs of Lemma 1, Theorem 1 and Proposition 1 are given in Section S2 of the Supplementary Material.
For comparison, we derive the expressions in this framework of the likelihood ratio test $W$, its Bartlett corrected test $W_{BC}$, and the two modifications $W^{*}$ and $W^{**}$ proposed by skovgaard:2001. The likelihood ratio test statistic (1) is
$W=n(\log|\hat{\Lambda}^{-1}_{0}|-\log|\hat{\Lambda}^{-1}|).$
Under $H_{\psi}$, its distribution is approximated by a $\chi^{2}_{d}$ with degrees of freedom d=p(g1)d=p(g-1) if and only if $p=o(n^{2/3})$ (He2020B). In this framework, the expectation $E(W)$ in the Bartlett corrected test (2) can be calculated exactly as done by He2020B. The $\chi^{2}_{d}$ approximation for the distribution of $W_{BC}$ holds if and only if $p=o(n^{4/5})$ (He2020B). The statistics $W^{*}$ and $W^{**}$ in (3) for hypothesis (11) can be computed explicitly, based on the formula (S1) in the Supplementary Material for the correction factor $\gamma$. The quantities required are $(\hat{\varphi}-\hat{\varphi}_{\psi})^{\top}(s-s_{\psi})={\text{tr}}(\hat{\Lambda}A)$, $\log(|{J}_{\varphi\varphi}(\hat{\varphi}_{\psi})|/|{J}_{\varphi\varphi}(\hat{\varphi})|)=(p+g+1)(\log|\hat{\Lambda}|-\log|\hat{\Lambda}_{0}|)$ and
$(s-s_{\psi})^{\top}J_{\varphi\varphi}(\hat{\varphi}_{\psi})^{-1}(s-s_{\psi})=$ $\left\{\sum_{i=1}^{g}n_{i}(\bar{y}_{i}-\bar{y})^{\top}\hat{\Lambda}_{0}(\bar{y}_{i}-\bar{y})\right\}$
$+\left\{\sum_{i=1}^{g}n_{i}(\bar{y}_{i}-\bar{y})^{\top}\right\}\frac{(\bar{y}^{\top}\hat{\Lambda}_{0}\bar{y})\hat{\Lambda}_{0}+\hat{\Lambda}_{0}\bar{y}\bar{y}^{\top}\hat{\Lambda}_{0}}{n}\left\{\sum_{i=1}^{g}n_{i}(\bar{y}_{i}-\bar{y})\right\}.$
3.2 Heteroscedastic one-way MANOVA
Suppose that $y_{ij}$ are independent observations from distributions $N_{p}(\mu_{i},\Lambda_{i}^{-1})$ for $i\in\{1,\dots,g\}$ and $j\in\{1,\dots,n_{i}\}$. We are again interested in testing the equality of the $g$ mean mean vectors $\mu_{i}$, i.e.
$H_{\psi}:\mu_{1}=\dots=\mu_{g}.$ (12)
In this framework, the constrained maximum likelihood estimates are denoted by $\hat{\mu}_{0i}=\hat{\mu}_{0}$ and $\hat{\Lambda}_{0i}^{-1}$, where $\hat{\Lambda}_{0i}^{-1}=\hat{\Lambda}_{i}^{-1}+(\bar{y}_{i}-\hat{\mu}_{0})(\bar{y}_{i}-\hat{\mu}_{0})^{\top}$. We compute the constrained maximum likelihood estimate $\hat{\mu}_{0}$ numerically by maximization of the profile log-likelihood $\ell_{P}(\mu)=-\sum_{i=1}^{g}(n_{i}/2)\log|\hat{\Lambda}_{i}^{-1}+$ $(\bar{y}_{i}-\mu)(\bar{y}_{i}-\mu)^{\top}|.$
To develop the directional test under the null hypothesis (12), following Fraser et al. (2016) we consider the parameterization $(\psi,\lambda)$ with parameter of interest $\psi=(\mu_{2}^{\top}-\mu_{1}^{\top},\dots,\mu_{g}^{\top}-\mu_{1}^{\top})^{\top}$ and nuisance parameter $\lambda=\{\mu_{1}^{\top},{\text{vech}}(\Lambda_{1})^{\top},$ $\ldots,{\text{vech}}(\Lambda_{g})^{\top}\}^{\top}$. This parameterization places nonlinear constraints on the canonical parameter $\varphi$. The tilted log-likelihood is $\ell(\varphi;t)=\sum_{i=1}^{g}\ell_{i}(\varphi_{i};t)$ with $\varphi_{i}=\{\mu_{i}^{\top}\Lambda_{i},\text{vech}(\Lambda_{i})^{\top}\}^{\top}$ and the $i$-th group’s contribution is
$\ell_{i}(\varphi_{i};t)=$ $\;{n_{i}}\mu_{i}^{\top}\Lambda_{i}\left\{t\bar{y}_{i}+(1-t)\hat{\mu}_{0}\right\}-\frac{1}{2}{\text{tr}\;}\left(\Lambda_{i}\left[y_{i}^{\top}y_{i}+(1-t)\left\{n_{i}\hat{\mu}_{0}(\bar{y}_{i}-\hat{\mu}_{0})^{\top}+n_{i}(\bar{y}_{i}-\hat{\mu}_{0})\hat{\mu}_{0}^{\top}\right\}\right]\right)$
$\;+\frac{n_{i}}{2}\log|\Lambda_{i}|-\frac{n_{i}}{2}\mu_{i}^{\top}\Lambda_{i}\mu_{i}.$
The expected value $s_{\psi}$ of the corresponding sufficient statistic $s$ under $H_{\psi}$ has components $n_{i}\hat{\mu}_{0}^{\top}-n_{i}\bar{y}_{i}^{\top},$ $\frac{n_{i}}{2}{\text{vech}}\{\hat{\mu}_{0}(\bar{y}_{i}-\hat{\mu}_{0})^{\top}+(\bar{y}_{i}-\hat{\mu}_{0})\hat{\mu}_{0}^{\top}\}^{\top}]$, $i\in\{1,\dots,g\}$. The maximum likelihood estimates along the line $s(t)=(1-t)s_{\psi}$ are $\hat{\mu}_{i}(t)=t\bar{y}_{i}+(1-t)\hat{\mu}_{0}$ and $\hat{\Lambda}_{i}^{-1}(t)=\hat{\Lambda}_{0i}^{-1}-t^{2}(\bar{y}_{i}-\hat{\mu}_{0})(\bar{y}_{i}-\hat{\mu}_{0})^{\top}$. The maximum likelihood estimate $\hat{\Lambda}_{i}^{-1}(t)$ exists for $t\leq t_{\text{sup}}=\mathop{\min}\limits_{1\leq i\leq g}\left[\{(\bar{y}_{i}-\hat{\mu}_{0})^{\top}\hat{\Lambda}_{0i}(\bar{y}_{i}-\hat{\mu}_{0})\}^{-1/2}\right]$, with $(\bar{y}_{i}-\hat{\mu}_{0})^{\top}\hat{\Lambda}_{0i}(\bar{y}_{i}-\hat{\mu}_{0})\neq 0$. Therefore, the saddlepoint approximation along the line $s(t)$ is
$h\{s(t);\psi\}=$ $\prod_{i=1}^{g}|\hat{\Lambda}_{i}^{-1}(t)|^{\frac{n_{i}-p-2}{2}}\left|\sum_{i=1}^{g}n_{i}\hat{\Lambda}_{0i}\left[I_{p}-t^{2}\left\{(p+1)I_{p}-{\text{tr}}(\hat{\Lambda}_{i}^{-1}\hat{\Lambda}_{0i})I_{p}-\hat{\Lambda}^{-1}_{i}\hat{\Lambda}_{0i}\right\}\right]\right|^{1/2}.$ (13)
In this case, the directional $p$-value is not expected to be exactly uniformly distributed under $H_{\psi}$, since there is no exact conditional density of the sufficient statistic that can be approximated by (13).
We also consider the likelihood ratio test and its modifications proposed by skovgaard:2001. The likelihood ratio statistic (1) takes the form
$W=\sum_{i=1}^{g}n_{i}\left(\log|\hat{\Lambda}_{0i}^{-1}|-\log|\hat{\Lambda}_{i}^{-1}|\right).$
The statistic $W$ has approximate $\chi^{2}_{d}$ null distribution with degrees of freedom d=p(g1)d=p(g-1), when $p$ is fixed. The expression of the correction factor $\gamma$, given in (S1), for hypothesis (12) in skovgaard:2001’s modifications (3) becomes
$\gamma=\frac{\left\{\sum_{i=1}^{g}n_{i}(\hat{\mu}_{0}-\bar{y}_{i})^{\top}\hat{\Lambda}_{0i}(\hat{\mu}_{0}-\bar{y}_{i})\right\}^{d/2}}{W^{d/2-1}\sum_{i=1}^{g}n_{i}(\hat{\mu}_{0}-\bar{y}_{i})^{\top}\hat{\Lambda}_{i}(\hat{\mu}_{0}-\bar{y}_{i})}\left\{\prod_{i=1}^{g}\frac{|\hat{\Lambda}_{i}|}{|\hat{\Lambda}_{0i}|}\right\}^{\frac{p+2}{2}}\left\{\frac{|\tilde{C}_{1}|}{|\sum_{i=1}^{g}n_{i}\hat{\Lambda}_{0i}|}\right\}^{1/2},$
where $\tilde{C}_{1}=\sum_{i=1}^{g}n_{i}\hat{\Lambda}_{0i}\left\{{\text{tr}\;}\left(\hat{\Lambda}_{i}^{-1}\hat{\Lambda}_{0i}-I_{p}\right)I_{p}+\hat{\Lambda}^{-1}_{i}\hat{\Lambda}_{0i}\right\}$.
If $g=2$, hypothesis (12) reduces to the multivariate Behrens-Fisher problem. Then, we can compare the directional test also with the procedure proposed by nel1986, which was shown to maintain a reasonable empirical type I error. The test by nel1986 is based on the quantity
$T^{*2}=(\bar{y}_{1}-\bar{y}_{2})^{\top}\left(\frac{S_{1}}{n_{1}}+\frac{S_{2}}{n_{2}}\right)^{-1}(\bar{y}_{1}-\bar{y}_{2}),$
where $S_{i}=\frac{n_{i}}{n_{i}-1}\hat{\Lambda}_{i}^{-1}$, $i=1,2$. The statistic $\frac{\upsilon-p+1}{p\upsilon}T^{*2}$ under $H_{\psi}$ has approximate $F$-distribution with degrees of freedom $(p,\upsilon-p+1)$, where
$\upsilon=\frac{{\text{tr}}\left\{\left(\frac{S_{1}}{n_{1}}+\frac{S_{2}}{n_{2}}\right)\left(\frac{S_{1}}{n_{1}}+\frac{S_{2}}{n_{2}}\right)\right\}+\left\{{\text{tr}}\left(\frac{S_{1}}{n_{1}}+\frac{S_{2}}{n_{2}}\right)\right\}^{2}}{\left.\left[{\text{tr}}\left\{\left(\frac{S_{1}}{n_{1}}\right)\left(\frac{S_{1}}{n_{1}}\right)\right\}+\left\{{\text{tr}}\left(\frac{S_{1}}{n_{1}}\right)\right\}^{2}\right]\right/{(n_{1}-1)}+\left[{\text{tr}}\left\{\left(\frac{S_{2}}{n_{2}}\right)\left.\left(\frac{S_{2}}{n_{2}}\right)\right\}+\left\{{\text{tr}}\left(\frac{S_{2}}{n_{2}}\right)\right\}^{2}\right]\right/{(n_{2}-1)}}.$
See rencher1998multivariate for more details.
All the different solutions will be evaluated by means of simulation studies in Section 4.2.
4 Simulation studies
4.1 Homoscedastic one-way MANOVA
The performance of the directional test for hypothesis (11) in the high dimensional multivariate normal framework is here assessed via Monte Carlo simulations based on $10,000$ replications. The exact directional test is compared with the chi-square approximations for $W$, $_{BC}, WW^{*}$, $W^{**}$, and with the normal approximation for the central limit theorem test proposed by He2020B, specifically for the high dimensional setting. The six tests are evaluated in terms of empirical size.
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},I_{p})$ under the null hypothesis. For each simulation experiment, we show results for $p=\kappa n_{i}$ with $\kappa\in\{0.05,0.1,0.3,$ 0.5,0.7,0.9}0.5,0.7,0.9\} and $n_{i}\in\{100,500\}$. Throughout, we set $g=3$ and $n_{1}=n_{2}=n_{3}$. In addition, we also consider some extreme settings with $\kappa\in\{1.0,1.5,2.0,2.5,2.9\}$ in which the chi-square approximations for $W$, $_{BC}, WW^{*}$, $W^{**}$ break down very fast (see Table 1).
Additional empirical results for the large sample size $n_{i}=1000$, and different values of $p$ and $g=30$ are reported in the Supplementary Material, which shows that the directional $p$-value maintains high accuracy.
Figure 2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left and right panels correspond to $n_{i}=100,500$, respectively.
The empirical size, i.e. the actual probability of type I error, at nominal level $\alpha=0.05$ based on the null distribution of the various statistics is reported in Figure S2. The directional $p$-value performs very well across all different values of $p$, confirming the exactness result in Theorem 1, while the central limit theorem test is less accurate when $p$ is small. The test based on $W$ breaks down in all settings. Instead, the Bartlett corrected test $W_{BC}$ proves accurate for moderate values of $p$, as seen in He2020B. However, the chi-square approximations for $_{BC}, WW^{*}$ and $W^{**}$ get unreliable as $p$ grows. Table 1 reports the empirical size in some extreme settings with $n_{i}\leq p\leq\sum_{i=1}^{g}n_{i}-g-1$. The results show that the chi-square approximations for $W$, $_{BC}, WW^{*}$ and $W^{**}$ do not work in such extreme settings, while the directional test is still accurate. The central limit theorem test, instead, becomes too liberal with large $p$.
Table 1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, $p=\kappa n_{i}$ and $n_{i}=100$, at nominal level $\alpha=0.05$.
$\kappa$ ($p$) DT CLT LRT BC Sko1 Sko2
1.0 (100) 0.048 0.053 0.677 0.066 0.064 0.050
1.5 (150) 0.048 0.055 0.991 0.133 0.131 0.073
2.0 (200) 0.046 0.052 1.000 0.401 0.402 0.145
2.5 (250) 0.052 0.057 1.000 0.967 0.966 0.567
2.9 (290) 0.048 0.060 1.000 1.000 1.000 0.999
4.2 Heteroscedastic one-way MANOVA
The performance of the directional test for hypothesis (12) in a moderate dimensional multivariate normal framework with $p=O(n_{i}^{\tau})$, $\tau\in(0,1)$ is here evaluated via Monte Carlo simulations based on 10,000 replications. The directional test is again compared with the chi-square approximations for the likelihood ratio test and its modifications proposed by skovgaard:2001. When $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$ by nel1986. The different testing approaches are evaluated in terms of empirical size. The results for $g=5,30$ are reported in Section S2.3 of the Supplementary Material and are in line with the ones discussed below.
We generate the data matrix $y_{i}$ as $n_{i}$ independent replications from a multivariate normal distributions $N_{p}(\mu_{i},\Lambda_{i}^{-1}),i\in\{1,\dots,g\}$. Under the null hypothesis $H_{\psi}$, we set $\mu_{i}=0_{p}$ and use an autoregressive structure for the covariance matrices, i.e. $\Lambda_{i}^{-1}=(\sigma_{jl})_{p\times p}=(\rho_{i}^{|j-l|})_{p\times p}$, with the $\rho_{i}$ chosen as an equally-spaced sequence from 0.1 to 0.9 of length $g$. In particular, when $g=2$, $\Lambda_{1}^{-1}=(0.1^{|j-l|})_{p\times p}$ and $\Lambda^{-1}_{2}=(0.9^{|j-l|})_{p\times p}$. We show results for $p=\lceil n_{i}\rceil^{\tau}$ with $\tau=j/24$, $j\in\{6,\dots,22\}$, $n_{i}\in\{100,500\}$ and $g=2$.
Figure 3: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (12) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left and right panels correspond to $n_{i}=100,500$, respectively.
Figure S10 reports the empirical size at the nominal level $\alpha=0.05$ under the null hypothesis for various statistics. The directional test is always more reliable than its competitors in terms of the empirical size, even if in this framework it is not exact. skovgaard:2001’s modifications are not as accurate when $p$ is large. Moreover, the likelihood ratio test and the Behrens-Fisher test break down quickly when $p=\lceil n_{i}\rceil^{\tau}$, even for moderately large values of $\tau$. Additional simulation results are shown in Table 2 for different structures of the covariance matrices: (I) identity matrix, $\Lambda_{i}^{-1}=I_{p}$; (II) compound symmetric matrix, $\Lambda^{-1}_{i}=(1-\rho_{i})I_{p}+\rho_{i}1_{p}1_{p}^{\top}$, with the same values of $\rho_{i}$ as in the autoregressive structure. Table 2 reports the empirical size at the nominal level $\alpha=0.05$ for $n_{i}=100$ and $p=\lceil n_{i}\rceil^{\tau}$, with covariance structures (I) and (II) and $g=2,30$. As expected, the likelihood ratio test performs in general very poorly, especially for larger values of $g$. skovgaard:2001’s modifications are much more accurate but not as much as the directional test, whose excellent performance is confirmed. When $g=2$, the approximate solution for the Behrens-Fisher test seems reliable in the identity covariance matrix case, but not under compound symmetry.
Table 2: Empirical size of the directional test (DT), likelihood ratio test (LRT), two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) (only for $g=2$) for hypothesis (12) at nominal level $\alpha=0.05$, with $n_{i}=100$ and $p=n_{i}^{\tau}$.
$\Lambda_{i}^{-1}$ $\tau$ ($p$) $g=2$ $g=30$
DT BF LRT Sko1 Sko2 DT LRT Sko1 Sko2
(I) 10/24 (7) 0.051 0.051 0.058 0.051 0.051 0.048 0.118 0.049 0.048
12/24 (10) 0.050 0.050 0.059 0.050 0.050 0.048 0.188 0.050 0.048
14/24 (15) 0.050 0.050 0.069 0.050 0.050 0.048 0.382 0.051 0.048
16/24 (22) 0.048 0.048 0.079 0.049 0.048 0.049 0.754 0.060 0.052
18/24 (32) 0.053 0.052 0.110 0.054 0.053 0.050 0.993 0.087 0.061
20/24 (47) 0.046 0.045 0.180 0.050 0.047 0.054 1.000 0.223 0.099
22/24 (69) 0.054 0.052 0.377 0.066 0.059 0.048 1.000 0.928 0.404
(II) 10/24 (7) 0.050 0.054 0.059 0.050 0.050 0.049 0.118 0.049 0.048
12/24 (10) 0.053 0.058 0.065 0.053 0.053 0.048 0.187 0.049 0.048
14/24 (15) 0.048 0.062 0.076 0.050 0.049 0.049 0.383 0.052 0.050
16/24 (22) 0.049 0.075 0.104 0.051 0.049 0.049 0.755 0.060 0.051
18/24 (32) 0.050 0.104 0.171 0.057 0.054 0.050 0.993 0.085 0.059
20/24 (47) 0.049 0.154 0.345 0.071 0.062 0.053 1.000 0.225 0.099
22/24 (69) 0.081 0.296 0.766 0.174 0.136 0.048 1.000 0.932 0.408
5 Applications
The dataset is taken from a study on how $p=5$ characteristics of the blood (red cell count, hemoglobin concentration, hematocrit, white cell count, and plasma ferritin concentration) vary with sport, body size, and sex of the athletes (telford1991sex). Our goal is to investigate whether the blood characteristics of athletes differ between endurance-related events and power-related events. The data can be found in the function ais of the R (Rcode) package DAAG (daag24). We performed several tests on this dataset to study the difference in blood parameters between athletes performing endurance-related and power-related sports. Typically, the conclusions drawn by the directional approach align with those from its best competitors. As an illustration, we report here some results involving two hypotheses: (i) testing the equality of blood characteristics between athletes competing in an endurance-related event (400-meter track event, $T_{400}$) and those in a power-related event (sprint, $T_{Sprint}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{T_{Sprint}}$; and (ii) testing the equality of blood characteristics for three groups of athletes that perform endurance-related sports ($Row$, SwimSwim, and $T_{400}$), i.e. $H_{0}:\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}$. Here we focus exclusively on data from male athletes as each blood variable is significantly different between genders (telford1991sex). We also consider the variable plasma ferritin concentration on the logarithmic scale. The sample sizes of $T_{400}$, $T_{Sprint}$, $Row$ and $Swim$ groups are respectively 18, 11, 15, 13. All the testing procedures under analysis can then be used to perform one-way MANOVA, and the corresponding $p$-values are shown in Table 3.
For hypothesis (i), the $p$-values in Table 3 of the directional test (homoscedastic: 0.060; heteroscedastic: 0.097), the Bartlett corrected test (homoscedastic: 0.059), the two Skovgaard’s modifications (homoscedastic: 0.078 and 0.083; heteroschedastic: 0.089 and 0.093) and the Behrens-Fisher test (heteroscedastic: 0.179) claim that the average blood characteristics levels for male athletes competing in $T_{400}$ and $T_{Sprint}$ events are not significantly different at a 5% level. Instead, the $p$-values of the central limit theorem test (homoscedastic: 0.036) and the likelihood ratio test (homoscedastic: 0.027; heteroscedastic: 0.034) both lead to a different conclusion. Note that the central limit theorem test is not reliable with small $p$ and $n_{i}$, while the likelihood ratio test is not valid if $\kappa$ and $\tau$ are large. To assess the validity of the various testing procedures in this setting, we can also examine the empirical size of the various tests. We generate $R=10,000$ samples with sizes $n_{T_{400}}=18$ and $n_{T_{Sprint}}=11$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{T_{Sprint}}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{T_{Sprint}}\bar{y}_{T_{Sprint}})/(n_{T_{400}}+n_{T_{Sprint}})$ where $\bar{y}_{T_{400}}$ and $\bar{y}_{T_{Sprint}}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}^{-1}_{0}$, the pooled sample covariance matrix for the homoscedastic case, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$ sample covariance matrices of each group, for the heteroscedastic case. The simulation results for hypothesis (i) reported in Table 4 show that the directional test is more reliable than its competitors, having empirical size closer to the nominal level $0.05$. The central limit theorem test and likelihood ratio test suffer from large bias. These results confirm the simulation outputs in Section S3.
Table 3: $p$-values of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
Hypothesis $g$ Homoscedastic Heteroscedastic
DT CLT LRT BC Sko1 Sko2 DT BF LRT Sko1 Sko2
(i) 2 0.060 0.036 0.027 0.059 0.078 0.083 0.097 0.179 0.034 0.089 0.093
(ii) 3 0.092 0.069 0.045 0.084 0.101 0.105 0.157 - 0.062 0.152 0.158
Table 4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) , two skovgaard:2001’s modifications (Sko1 and Sko2), and Behrens-Fisher test (BF) for testing whether the blood characteristics of athletes differ between endurance-related and power-related events at nominal level $\alpha=0.05$, with $n_{i}\in\{18,11,15,13\}$ and $p=5$.
Hypothesis $g$ Homoscedastic Heteroscedastic
DT CLT LRT BC Sko1 Sko2 DT BF LRT Sko1 Sko2
(i) 2 0.049 0.070 0.092 0.050 0.038 0.034 0.053 0.017 0.111 0.055 0.053
(ii) 3 0.045 0.062 0.087 0.046 0.039 0.038 0.049 - 0.125 0.050 0.047
For hypothesis (ii), all the tests used under heteroscedasticity in Table 3 give $p$-values directional: 0.097; likelihood ratio: 0.062; Skovgaard’s: 0.152 and 0.158) concluding that differences in mean blood characteristics among endurance-related events ($T_{400}$, $Row$ and $Swim$) for male athletes are not statistically significant at a 5% level. In the homoscedastic case, only the likelihood ratio test yields a significant $p$-value (0.045). Conversely, the other test statistics have $p$-values (directional: 0.092; central limit theorem: 0.069; Bartlett corrected: 0.084; Skovgaard’s: 0.101 and 0.105) that agree with the hypothesis of no significant difference among the three groups. Even in this case, a simulation study can be performed in order to verify the reliability of the considered methods. Under the null hypothesis, we generate $10,000$ samples with sizes $n_{T_{400}}=18$, $n_{Row}=15$ and $n_{Swim}=13$ from a normal distribution with mean vector $\mu_{T_{400}}=\mu_{Row}=\mu_{Swim}=(n_{T_{400}}\bar{y}_{T_{400}}+n_{Row}\bar{y}_{Row}+n_{Swim}\bar{y}_{Swim})/(n_{T_{400}}+n_{Row}+n_{Swim})$ where $\bar{y}_{T_{400}}$, $\bar{y}_{Row}$ and $\bar{y}_{Swim}$ are the group sample means, and covariance matrix $\Lambda^{-1}=\hat{\Lambda}_{0}^{-1}$, the pooled sample covariance matrix under the homoscedasticity, and covariance matrices $\Lambda^{-1}_{i}=\hat{\Lambda}_{i}^{-1}$, sample covariance matrices under the heteroscedasticity. Simulation results are reported in Table 4. In the homoscedastic case, the two Skovgaard’s modifications are conservative. The central limit theorem test and the likelihood test, on the other hand, have slightly inflated empirical size. Hence, the $p$-values of these four methods are not reliable in this scenario. In contrast, the directional test and the Bartlett corrected test well control the type I error. In the heteroscedastic case, the likelihood ratio test is overly inflated, whereas the remaining procedures, including the directional test, exhibit a good performance.
6 Conclusions
In this paper, we have developed the directional test for one-way MANOVA problems when the data dimension is comparable with the sample size. The directional $p$-value has been proved exactly uniformly distributed in a high dimensional setting, provided that $\sum_{i=1}^{g}n_{i}\geq p+g+1$, when testing the equality of normal mean vectors with identical covariance matrix. This finding is supported by our numerical studies. In addition, simulations in moderate dimensional scenarios in the heteroscedastic one-way MANOVA framework indicate that the directional test outperforms its competitors in terms of empirical null distribution, even in the more general setting with different covariance matrices. In general, all the considered approaches rely on the assumed multivariate normal model and are not expected to be reliable if that is misspecified. Yet, our numerical results show that these deviations from normality do not affect much the behaviour of the various tests and identify the directional test as the preferable solution even under these model misspecifications. While formal conditions for the validity of the various methods in this scenario could perhaps be developed using recent results on the accuracy of the saddlepoint approximation in moderate dimensional regimes (tang2021laplace), we believe further work is needed in this area to refine these methods under more complex settings. It is also worth mentioning that, given the inconclusive results of the power analysis of likelihood-based tests for high dimensional normal distributions included in huang2021directional, here we have chosen not to compare the performance of the tests under a few arbitrary alternative hypotheses. Further investigations have showed that all the likelihood-based solutions examined are reasonably robust to the misspecification of the assumed multivariate normal model, although different and/or more substantial deviations from normality could be considered in future research for a more complete picture.
Supplementary Material
Supplementary Material provides auxiliary computational details, formal proofs, and extensive additional simulation studies covering both homoscedastic and heteroscedastic high and moderate dimensional one-way MANOVA, along with an applicaiton results from breast cancer microarray data. The R code to reproduce all numerical results in the paper are available at https://github.com/stat-cz/DirTestMANOVA.
Supplementary material to Directional testing for one-way MANOVA in divergent dimensions
S1 Computational results
S1.1 Correction factor $\gamma$ in skovgaard:2001’s modifications
The general expression for $\gamma$ appearing in skovgaard:2001’s modified versions of W (3) in the main text can be written as
$\gamma=\frac{\left\{(s-s_{\psi})^{\top}J_{\varphi\varphi}(\hat{\varphi}_{\psi})^{-1}(s-s_{\psi})\right\}^{d/2}}{W^{d/2-1}(\hat{\varphi}-\hat{\varphi}_{\psi})^{\top}(s-s_{\psi})}\left\{\frac{|{J}_{\varphi\varphi}(\hat{\varphi}_{\psi})|}{|{J}_{\varphi\varphi}(\hat{\varphi})|}\right\}^{1/2}\left\{\frac{|{J}_{\lambda\lambda}(\hat{\varphi}_{\psi})|}{|{I}_{\lambda\lambda}(\hat{\varphi}_{\psi})|}\right\}^{1/2},$ (S14)
where $J_{\lambda\lambda}$ and $I_{\lambda\lambda}$ are the blocks of observed and expected information matrices, respectively, relative to the nuisance parameter $\lambda$. For calculating the $p$-value, the quantity (S14) is evaluated at $s=s^{0}=0_{q}$, corresponding to the observed data point $y^{0}$. In particular, if $\lambda$ is a component or a linear function of the canonical parameter, as in Section 3.1 in the main text, the last factor in (S14) equals 1.
S1.2 Observed information matrix
When testing hypothesis (8) in the main text, the saddlepoint approximation (salvan:1997, Section 10) and the correction factor $\gamma$ in (3) require the computation of the observed information matrix $J_{\varphi\varphi}(\varphi)$. This has the block form
$J_{\varphi\varphi}(\varphi)$ $=\left(\begin{array}[]{cc}J_{\xi\xi}&J_{\xi\text{vech}(\Lambda)}\\ J_{\text{vech}(\Lambda)\xi}&J_{\text{vech}(\Lambda)\text{vech}(\Lambda)}\end{array}\right),$ (S17)
$\text{with}\;J_{\xi\xi}=\left(\begin{array}[]{cccc}n_{1}\Lambda^{-1}&0_{p\times p}&\cdot&0_{p\times p}\\ 0_{p\times p}&n_{2}\Lambda^{-1}&\cdot&0_{p\times p}\\ \vdots&\vdots&\vdots&\vdots\\ 0_{p\times p}&0_{p\times p}&\cdot&n_{g}\Lambda^{-1}\end{array}\right),\;J_{\text{vech}(\Lambda)\xi}^{\top}=J_{\xi\text{vech}(\Lambda)}=\left(\begin{array}[]{c}-n_{1}(\xi_{1}^{\top}\Lambda^{-1}\otimes\Lambda^{-1})D_{p}\\ -n_{2}(\xi_{2}^{\top}\Lambda^{-1}\otimes\Lambda^{-1})D_{p}\\ \vdots\\ -n_{k}(\xi_{g}^{\top}\Lambda^{-1}\otimes\Lambda^{-1})D_{p}\end{array}\right)$ (S26)
and $J_{\text{vech}(\Lambda)\text{vech}(\Lambda)}=\sum_{i=1}^{g}\frac{n_{i}}{2}D_{p}^{\top}\Lambda^{-1}(I_{p}+2\xi_{i}\xi_{i}^{\top}\Lambda^{-1})\otimes\Lambda^{-1}D_{p},$ where $\otimes$ denotes the Kronecker product (see, for instance, Lauritzen:1996). Then, the determinant of $J_{\varphi\varphi}(\varphi)$ can be computed as
$|J_{\varphi\varphi}(\varphi)|=|J_{\xi\xi}||C_{2}|,$
where $C_{2}=J_{\text{vech}(\Lambda)\text{vech}(\Lambda)}-J_{\text{vech}(\Lambda)\xi}J_{\xi\xi}^{-1}J_{\xi\text{vech}(\Lambda)}=\sum_{i=1}^{g}\frac{n_{i}}{2}D_{p}^{\top}\Lambda^{-1}(I_{p}+2\xi_{i}\xi_{i}^{\top}\Lambda^{-1})\otimes\Lambda^{-1}D_{p}-\sum_{i=1}^{g}n_{i}D_{p}^{\top}\Lambda^{-1}$ $\xi_{i}\xi_{i}^{\top}$ $\Lambda^{-1}\otimes\Lambda^{-1}D_{p}$ $\frac{n}{2}D_{p}^{\top}\left(\Lambda^{-1}\otimes\Lambda^{-1}\right)D_{p}$.
After some algebra, we get
$|J_{\varphi\varphi}(\varphi)|=\left(\prod_{i=1}^{g}n_{i}^{p}\right)|\Lambda^{-1}|^{g}n^{\frac{p(p+1)}{2}}2^{-p}|\Lambda^{-1}|^{p+1}\propto|\Lambda^{-1}|^{p+g+1}.$
S2 Proofs
S2.1 Proof of Lemma 1
Recall that $\hat{\Lambda}^{-1}(t)=\hat{\Lambda}_{0}^{-1}-t^{2}A/n=\hat{\Lambda}^{-1}+(1-t^{2})A/n$ with $\hat{\Lambda}_{0}^{-1}=\hat{\Lambda}^{-1}+A/n$, where $A=\sum_{i=1}^{g}n_{i}(\bar{y}_{i}-\bar{y})(\bar{y}_{i}-\bar{y})^{\top}$. If $t\in[0,1]$ the result is straightforward, since, for all $x\in\mathbb{R}^{p}$, $x^{\top}\hat{\Lambda}^{-1}(t)x=x^{\top}\hat{\Lambda}^{-1}x/n+(1-t^{2})x^{\top}Ax/n>0$. Let us focus on the case $t>1$. We rewrite the estimator $\hat{\Lambda}^{-1}(t)$ as
$\hat{\Lambda}^{-1}(t)=B_{0}^{\top}\left\{I_{p}-t^{2}(B_{0}^{\top})^{-1}(A/n)B_{0}^{-1}\right\}B_{0},$
with $B_{0}$ such that $\hat{\Lambda}_{0}^{-1}=B_{0}^{\top}B_{0}$. According to the eigen decomposition, the matrix $(B_{0}^{\top})^{-1}(A/n)B_{0}^{-1}=PQP^{\top}$ with an orthogonal matrix $P$ whose columns are eigenvectors of $(B_{0}^{\top})^{-1}(A/n)B_{0}^{-1}$ and a diagonal matrix $Q$ whose diagonal elements are the eigenvalues of $(B_{0}^{\top})^{-1}(A/n)B_{0}^{-1}$. Then, we have $\hat{\Lambda}^{-1}(t)=B_{0}^{\top}P\{I_{p}-t^{2}Q\}P^{\top}B_{0}$. Therefore, checking that $\hat{\Lambda}^{-1}(t)$ is positive definite is equivalent to checking that $I_{p}-t^{2}Q$ is positive definite. Indeed, for all $x\in\mathbb{R}^{p},x\neq 0$, then
$x^{\top}\hat{\Lambda}^{-1}x$ $=$ $x^{\top}B^{\top}_{0}P\{I_{p}-t^{2}Q\}P^{\top}B_{0}x$
$=$ $\tilde{x}^{\top}\{I_{p}-t^{2}Q\}\tilde{x}>0,$
where $\tilde{x}=P^{\top}B_{0}x$, with $\tilde{x}\neq 0$ if $x\neq 0$.
Next, the positive definiteness of $I_{p}-t^{2}Q$ should be proved. It is equivalent to checking that all elements of the diagonal matrix $I_{p}-t^{2}Q=\text{diag}(1-t^{2}\nu_{l})$ are positive, where $\nu_{l},l\in\{1,\dots,p\}$, are the eigenvalues of the matrix $(B_{0}^{\top})^{-1}(A/n)B_{0}^{-1}$. We now need to find out the largest $t$ such that $1-t^{2}\nu_{l}>0$, $l\in\{1,\dots,p\}$. Then, the largest value of $t$ for which $\hat{\Lambda}^{-1}(t)$ is positive definite equals $t_{\text{sup}}=\sqrt{1/\nu_{(p)}}$, where $\nu_{p}$ is the maximum eigenvalue of the matrix $(B_{0}^{-1})^{\top}AB_{0}^{-1}/n$. Therefore, $\hat{\Lambda}^{-1}(t)$ is positive definite in $t\in[0,\sqrt{1/\nu_{(p)}}]$.
S2.2 Proof of Theorem 1
Suppose $y_{i}=[y_{i1}\cdots y_{in_{i}}]^{\top}$, with $y_{ij}\sim N_{p}(\mu_{i},\Lambda^{-1})$, $i\in\{1,\dots,g\}$ and $j\in\{1,\dots,n_{i}\}$. The log-likelihood for the canonical parameter $\varphi$ is
$\ell(\varphi;s)=\sum_{i=1}^{g}-\frac{n_{i}}{2}\log|\Lambda^{-1}|-\frac{1}{2}{\text{tr}}(\Lambda B_{i})-\frac{n_{i}}{2}(\bar{y}_{i}-\mu_{i})^{\top}\Lambda(\bar{y}_{i}-\mu_{i}),$ (S27)
[…]
$h(s;\psi)$ $=$ $\left[\prod_{i=1}^{g}c_{i1}|\hat{\Lambda}_{0}^{-1}|^{-\frac{1}{2}}\exp\left\{-\frac{n_{i}}{2}(\bar{y}_{i}-\hat{\mu}_{\psi})^{\top}\hat{\Lambda}_{0}(\bar{y}_{i}-\hat{\mu}_{\psi})\right\}\right]$ (S28)
$\;\;\times c_{2}|\hat{\Lambda}_{0}^{-1}|^{\frac{n-1}{2}}\exp\left\{-\frac{n}{2}{\text{tr}}(\hat{\Lambda}_{0}\hat{\Lambda}^{-1})\right\}|\hat{\Lambda}^{-1}|^{\frac{n-p-g-1}{2}}.$
Expression (S28) equals the exact joint distribution of $\bar{y}_{1},\dots,\bar{y}_{g}$ and $\hat{\Lambda}^{-1}$ if $c_{i1}=(2\pi)^{-p/2}$ and $c_{2}=\left(\frac{n}{2}\right)^{p(n-1)/2}$ $\Gamma_{p}\left(\frac{n-g}{2}\right)^{-1}$ and with fixed values of $\hat{\mu}_{\psi}$ and $\hat{\Lambda}_{0}^{-1}$. Indeed, the constrained maximum likelihood estimates are fixed and equal to their observed values when we consider the saddlepoint approximation along the line $s(t)$ under $H_{\psi}$. Moreover, we have the unconstrained maximum likelihood estimates $\hat{\mu}(t)=t\bar{y}_{i}+(1-t)\bar{y}$ and $\hat{\Lambda}^{-1}(t)=\hat{\Lambda}_{0}^{-1}-t^{2}A/n$. Then, the saddlepoint approximation to the conditional distribution of $s(t)$ under $H_{\psi}$ follows from (S28) and is equal to
$h\{s(t);\psi\}$ $=$ $\prod_{i=1}^{g}c_{i1}|\hat{\Lambda}_{0}^{-1}|^{-\frac{1}{2}}\exp\left[-\frac{n_{i}}{2}\{\hat{\mu}_{i}(t)-\hat{\mu}_{\psi}\}^{\top}\hat{\Lambda}_{0}\{\hat{\mu}_{i}(t)-\hat{\mu}_{\psi}\}\right]$
$\;\;\times c_{2}|\hat{\Lambda}_{0}^{-1}|^{\frac{n-1}{2}}\exp\left[-\frac{n}{2}{\text{tr}}\{\hat{\Lambda}_{0}\hat{\Lambda}(t)^{-1}\}\right]|\hat{\Lambda}(t)^{-1}|^{\frac{n-p-g-1}{2}}$
$\propto$ $\prod_{i=1}^{g}\exp\left[-\frac{n_{i}}{2}\{t\bar{y}_{i}+(1-t)\bar{y}-\bar{y}\}^{\top}\hat{\Lambda}_{0}\{t\bar{y}_{i}+(1-t)\bar{y}-\bar{y}\}\right]$
$\;\;\times\exp\left[-\frac{n}{2}{\text{tr}}\{\hat{\Lambda}_{0}(\hat{\Lambda}_{0}^{-1}-t^{2}A/n)\}\right]|\hat{\Lambda}^{-1}(t)|^{\frac{n-p-g-1}{2}}$
$\propto$ $\exp\left\{-\frac{1}{2}t^{2}\sum_{i=1}^{g}(\bar{y}_{i}-\bar{y})^{\top}\hat{\Lambda}_{0}(\bar{y}_{i}-\bar{y})\right\}\exp\left\{\frac{1}{2}t^{2}{\text{tr}}(\hat{\Lambda}_{0}A)\right\}|\hat{\Lambda}^{-1}(t)|^{\frac{n-p-g-1}{2}}$
$\propto$ $|\hat{\Lambda}^{-1}(t)|^{\frac{n-p-g-1}{2}}.$
[…]
S2.3 Proof of Proposition 1
First, we express the estimate $\hat{\Lambda}^{-1}(t)=\hat{\Lambda}_{0}^{-1}-t^{2}A/n$, where $n=n_{1}+n_{2}$, as
$\hat{\Lambda}^{-1}(t)=$ $\hat{\Lambda}_{0}^{-1}-t^{2}\sum_{i=1}^{2}n_{i}(\bar{y}_{i}-\bar{y})(\bar{y}_{i}-\bar{y})^{\top}/n,$
where $\bar{y}=(n_{1}\bar{y}_{1}+n_{2}\bar{y}_{2})/n$. Moreover, we have
$\sum_{i=1}^{2}n_{i}(\bar{y}_{i}-\bar{y})(\bar{y}_{i}-\bar{y})^{\top}=$ $\;n_{1}(\bar{y}_{1}-\bar{y})(\bar{y}_{1}-\bar{y})^{\top}+n_{2}(\bar{y}_{2}-\bar{y})(\bar{y}_{2}-\bar{y})^{\top}$
$=$ $\;\frac{n_{1}n_{2}(\bar{y}_{1}-\bar{y}_{2})(\bar{y}_{1}-\bar{y}_{2})^{\top}}{n},$
since
$n_{1}\bar{y}_{1}\bar{y}_{1}^{\top}+n_{2}\bar{y}_{2}\bar{y}_{2}^{\top}=$ $\frac{n_{1}^{2}\bar{y}_{1}\bar{y}_{1}^{\top}+n_{1}n_{2}\bar{y}_{1}\bar{y}_{1}^{\top}+n_{1}n_{2}\bar{y}_{2}\bar{y}_{2}^{\top}+n_{2}^{2}\bar{y}_{2}\bar{y}_{2}^{\top}}{n},$
$n\bar{y}\bar{y}^{\top}=$ $\frac{n_{1}^{2}\bar{y}_{1}\bar{y}_{1}^{\top}+n_{1}n_{2}\bar{y}_{1}\bar{y}_{2}^{\top}+n_{1}n_{2}\bar{y}_{2}\bar{y}_{1}^{\top}+n_{2}^{2}\bar{y}_{2}\bar{y}_{2}^{\top}}{n}.$
Then
$\hat{\Lambda}^{-1}(t)=$ $\hat{\Lambda}_{0}^{-1}-t^{2}\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})(\bar{y}_{1}-\bar{y}_{2})^{\top},$
$|\hat{\Lambda}^{-1}(t)|=$ $|\hat{\Lambda}_{0}^{-1}|\left\{1-t^{2}\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}_{0}(\bar{y}_{1}-\bar{y}_{2})\right\}.$
The integrand function in the directional $p$-value (7) along the line $s(t)$ can then be simplified to
$g(t)=t^{d-1}|\hat{\Lambda}^{-1}(t)|^{\frac{n-p-3}{2}}\propto t^{d-1}\left\{1-t^{2}\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}_{0}(\bar{y}_{1}-\bar{y}_{2})\right\}^{\frac{n-p-3}{2}}.$
Thus the directional $p$-value under the null hypothesis $H_{\psi}$ takes the form
$p(\psi)=\frac{\int_{1}^{t_{\text{sup}}}t^{d-1}\left\{1-t^{2}\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}_{0}(\bar{y}_{1}-\bar{y}_{2})\right\}^{\frac{n-p-3}{2}}\text{d}t}{\int_{0}^{t_{\text{sup}}}t^{d-1}\left\{1-t^{2}\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}_{0}(\bar{y}_{1}-\bar{y}_{2})\right\}^{\frac{n-p-3}{2}}\text{d}t},$
where $t_{\text{sup}}=\left\{\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}_{0}(\bar{y}_{1}-\bar{y}_{2})\right\}^{-1/2}$. In order to make the notation more compact, we define $C=\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}_{0}(\bar{y}_{1}-\bar{y}_{2})$, so that $t_{\text{sup}}=C^{-1/2}$. We can now rewrite the directional $p$-value as
$p(\psi)=\frac{\int_{1}^{C^{-1/2}}t^{d-1}\left\{1-t^{2}C\right\}^{\frac{n-p-3}{2}}\text{d}t}{\int_{0}^{C^{-1/2}}t^{d-1}\left\{1-t^{2}C\right\}^{\frac{n-p-3}{2}}\text{d}t}.$ (S29)
Since the Hotelling’s $T^{2}$ statistic has Hotelling’s $T^{2}$ distribution, i.e., $T^{2}\sim T^{2}(p,n-p-1)$ with degrees of freedom $df_{1}=p$ and $df_{2}=n-p-1$, we change variable from $t$ to $\left\{C\left(x{df_{2}}/{df_{1}}+1\right)\right\}^{-1/2}$.
The following steps are used to compute the numerator in the directional $p$-value (S29).
Step 1. Change of the integration interval:
$1\leq t\leq C^{-1/2}\Leftrightarrow 1\leq\left\{C\left(\frac{df_{2}}{df_{1}}x+1\right)\right\}^{-1/2}\leq C^{-1/2}$
$\Leftrightarrow C\leq\left(\frac{df_{2}}{df_{1}}x+1\right)^{-1}\leq 1\Leftrightarrow 1\leq\frac{df_{2}}{df_{1}}x+1\leq C^{-1}$
$\Leftrightarrow 0\leq\frac{df_{2}}{df_{1}}x\leq C^{-1}-1\Leftrightarrow 0\leq x\leq\frac{df_{1}}{df_{2}}\frac{1-C}{C}$
hence the new integral is on $\left[0,\frac{df_{1}}{df_{2}}\frac{1-C}{C}\right]$.
Step 2. Change of variable from $t$ to $x$:
$\int_{1}^{C^{-1/2}}t^{d-1}\left\{1-t^{2}C\right\}^{\frac{n-p-3}{2}}\text{d}t$
$=$ $\int_{0}^{\frac{df_{1}}{df_{2}}\frac{1-C}{C}}\left\{C\left(\frac{df_{2}}{df_{1}}x+1\right)\right\}^{-\frac{d-1}{2}}\left\{1-C^{-1}\left(\frac{df_{2}}{df_{1}}x+1\right)^{-1}C\right\}^{\frac{n-p-3}{2}}\text{d}\left\{C\left(\frac{df_{2}}{df_{1}}x+1\right)\right\}^{-\frac{1}{2}},$
where d=p(k1)=pd=p(k-1)=p. Then we get
${\text{d}\left\{C\left(\frac{df_{2}}{df_{1}}x+1\right)\right\}^{-\frac{1}{2}}}=C^{-1/2}(-\frac{1}{2})\left(\frac{df_{2}}{df_{1}}x+1\right)^{-\frac{3}{2}}\frac{df_{2}}{df_{1}}{\text{d}x},$
and
$\int_{1}^{C^{-1/2}}t^{d-1}\left\{1-t^{2}C\right\}^{\frac{n-p-3}{2}}\text{d}t$
$=$ $c\int_{0}^{\frac{df_{1}}{df_{2}}\frac{1-C}{C}}\left(\frac{df_{2}}{df_{1}}x+1\right)^{-\frac{p-1}{2}}\left\{1-\left(\frac{df_{2}}{df_{1}}x+1\right)^{-1}\right\}^{\frac{n-p-3}{2}}\left(\frac{df_{2}}{df_{1}}x+1\right)^{-\frac{3}{2}}\text{d}x$
$=$ $c\int_{0}^{\frac{df_{1}}{df_{2}}\frac{1-C}{C}}\left(\frac{df_{2}}{df_{1}}x+1\right)^{-\frac{p-1+3}{2}}\left\{\frac{\frac{df_{2}}{df_{1}}x+1-1}{\frac{df_{2}}{df_{1}}x+1}\right\}^{\frac{n-p-3}{2}}\text{d}x$
$=$ $c\int_{0}^{\frac{df_{1}}{df_{2}}\frac{1-C}{C}}\left(\frac{df_{2}}{df_{1}}x+1\right)^{-\frac{n-p-1+p}{2}}\left(\frac{df_{2}}{df_{1}}x\right)^{\frac{n-p-1}{2}-1}\text{d}x$
$=$ $c\int_{0}^{\frac{df_{1}}{df_{2}}\frac{1-C}{C}}\left(\frac{df_{2}}{df_{1}}x+1\right)^{-\frac{df_{2}+df_{1}}{2}}\left(\frac{df_{2}}{df_{1}}x\right)^{\frac{df_{2}}{2}-1}\text{d}x$
$=$ $c^{\prime}F\left(\frac{df_{1}}{df_{2}}\frac{1-C}{C},df_{2},df_{1}\right),$
where $F(X,df_{2},df_{1})$ is the cumulative distribution function of a random variable $X$ following a $F$-distribution with degrees of freedom $df_{2}$ and $df_{1}$. Since if $X\sim F(df_{2},df_{1})$, then $X^{-1}\sim F(df_{1},df_{2})$, we can express the directional $p$-value as
$p(\psi)=1-F\left(\frac{df_{2}}{df_{1}}\frac{C}{1-C},df_{1},df_{2}\right).$
According to the Sherman-Morrison formula (see also sartori:2019Ftest, Section 4.2), we have that
$C=\left\{\left.{n_{1}n_{2}/n}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}(\bar{y}_{1}-\bar{y}_{2})\right\}\right/\left\{1+{n_{1}n_{2}/{n}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}(\bar{y}_{1}-\bar{y}_{2})\right\}$ and
$\frac{df_{2}}{df_{1}}\frac{C}{1-C}=$ $\frac{n-p-1}{p}\cdot\frac{n_{1}n_{2}}{n^{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}\hat{\Lambda}(\bar{y}_{1}-\bar{y}_{2})$
$=$ $\frac{n-p-1}{p}\cdot\frac{T^{2}}{n_{1}+n_{2}-2},$
with $T^{2}=\left.{n_{1}n_{2}}(\bar{y}_{1}-\bar{y}_{2})^{\top}S^{-1}(\bar{y}_{1}-\bar{y}_{2})\right/{n}$ and $S=\left.{n}\hat{\Lambda}^{-1}\right/{(n_{1}-n_{2}-2)}$. Since $\left.{(n-p-1)}{T^{2}}\right/\{p(n_{1}+n_{2}-2)\}\sim F(p,n-p-1)$, the directional test is identical to the Hotelling $T^{2}$ test.
S3 Simulation studies for Homoscedastic one-way MANOVA
This section reports additional empirical result for homoscedastic one-way MANOVA in the multivariate normal framework. We compare the performance of exact directional test (DT) with other five approximate approaches: the central limite theorem test (CLT), log-likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2). The six tests are evaluated in terms of empirical size. The simulation results are computed via Monte Carlo simulation based on 10,000 replications.
Figure S1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to $n_{i}=1000$, respectively.
S3.1 Empirical results for high dimentional setup
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},I_{p})$ under the null hypothesis. For each simulation experiment, we show results for $p=\kappa n_{i}$ with $\kappa\in\{0.05,0.1,0.3,$ 0.5,0.7,0.9}0.5,0.7,0.9\} and $n_{i}=1000$. Throughout, we set $g=3$ and $n_{1}=n_{2}=n_{3}$. Figure S1 shows the empirical size of the directional test, central limit theorem test, likelihood ratio test, Bartlett corrected test and two skovgaard:2001’s modifications for hypothesis (8) with $g=3$, at nominal level $\alpha=0.05$. The directional $p$-value performs better than the other. The log-likelihood ratio test is disastrous in the large sample size setting with $n_{i}=1000$.
Table S1: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $n_{i}=100,500,1000$, at nominal level $\alpha=0.05$.
$n_{i}$ $\tau(p)$ DT CLT LRT BC Sko1 Sko2
100 0.250 (3) 0.054 0.072 0.057 0.054 0.053 0.053
0.333 (4) 0.047 0.063 0.051 0.047 0.046 0.046
0.417 (6) 0.050 0.066 0.056 0.050 0.049 0.049
0.500 (10) 0.048 0.060 0.057 0.048 0.046 0.046
0.580 (14) 0.052 0.064 0.067 0.052 0.050 0.050
0.667 (21) 0.050 0.061 0.080 0.049 0.048 0.047
0.750 (31) 0.048 0.059 0.104 0.049 0.046 0.046
0.833 (46) 0.049 0.057 0.161 0.051 0.049 0.048
0.917 (68) 0.050 0.057 0.311 0.056 0.052 0.048
500 0.250 (4) 0.051 0.068 0.052 0.051 0.051 0.051
0.333 (7) 0.053 0.070 0.055 0.053 0.053 0.053
0.417 (13) 0.051 0.063 0.054 0.051 0.051 0.051
0.500 (22) 0.050 0.061 0.056 0.050 0.050 0.050
0.580 (37) 0.048 0.056 0.058 0.048 0.048 0.048
0.667 (62) 0.055 0.063 0.078 0.055 0.054 0.054
0.750 (105) 0.050 0.054 0.105 0.051 0.050 0.049
0.833 (177) 0.054 0.058 0.222 0.056 0.055 0.054
0.917 (297) 0.047 0.050 0.609 0.057 0.056 0.049
1000 0.250 (5) 0.054 0.070 0.055 0.054 0.054 0.054
0.333 (9) 0.052 0.066 0.053 0.053 0.052 0.052
0.417 (17) 0.049 0.061 0.050 0.049 0.049 0.049
0.500 (31) 0.054 0.063 0.058 0.054 0.054 0.054
0.580 (56) 0.052 0.059 0.061 0.052 0.052 0.052
0.667 (99) 0.049 0.055 0.072 0.049 0.049 0.049
0.750 (177) 0.053 0.056 0.116 0.053 0.053 0.052
0.833 (316) 0.048 0.050 0.261 0.049 0.049 0.048
0.917 (562) 0.053 0.055 0.794 0.064 0.065 0.056
S3.2 Empirical results for moderate dimentional setup
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},I_{p})$ under the null hypothesis. For each simulation experiment, we show results for $p=\lfloor n_{i}^{\tau}\rfloor$ with $\tau=j/24$, $j\in\{6,8,\cdots,22\}$ and $n_{i}\in\{100,500,1000\}$. Throughout, we set $g=3$ and $n_{1}=n_{2}=n_{3}$.
Figure S2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to $n_{i}=100,500,1000$, respectively ($g=3$).
S3.3 Empirical results for large number $g$ of groups
Table S2: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$, $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$
$\tau(p)$ DT CLT LRT BC Sko1 Sko2
0.250 (3) 0.053 0.060 0.057 0.053 0.053 0.053
0.333 (4) 0.050 0.057 0.055 0.050 0.050 0.050
0.417 (6) 0.050 0.056 0.057 0.050 0.050 0.050
0.500 (10) 0.049 0.053 0.059 0.049 0.049 0.049
0.580 (14) 0.051 0.056 0.064 0.052 0.051 0.051
0.667 (21) 0.050 0.054 0.071 0.050 0.050 0.050
0.750 (31) 0.052 0.054 0.080 0.052 0.052 0.052
0.833 (46) 0.048 0.051 0.095 0.049 0.048 0.048
0.917 (68) 0.054 0.056 0.137 0.054 0.054 0.053
Figure S3: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S3.4 Empirical results for the setup of He2020B
In each Monte Carlo experiment, we show results for $p=\lfloor n^{\tau}\rfloor$ with $n=\sum_{i=1}^{g}n_{i}$. Under the null hypothesis, we set $g=3$, $n_{1}=n_{2}=n_{3}$, $\tau=j/24$ with $j\in\{6,8,\cdots,22\}$ and $n_{i}=100$.
Table S3: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$.
$\tau$ ($p$) DT CLT LRT BC Sko1 Sko2
0.250 (4) 0.048 0.064 0.052 0.047 0.046 0.046
0.333 (6) 0.049 0.066 0.056 0.049 0.047 0.047
0.417 (10) 0.050 0.064 0.061 0.050 0.049 0.049
0.500 (17) 0.053 0.065 0.074 0.053 0.051 0.051
0.583 (27) 0.046 0.056 0.088 0.047 0.045 0.045
0.667 (44) 0.051 0.060 0.157 0.053 0.050 0.049
0.750 (72) 0.048 0.054 0.347 0.054 0.052 0.047
0.833 (115) 0.050 0.056 0.832 0.078 0.077 0.057
0.917 (186) 0.044 0.049 1.000 0.280 0.278 0.106
Figure S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}100=300$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S4: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$.
$\tau$ ($p$) DT CLT LRT BC Sko1 Sko2
0.250 (6) 0.052 0.068 0.053 0.052 0.052 0.052
0.333 (11) 0.051 0.063 0.053 0.051 0.050 0.050
0.417 (21) 0.050 0.059 0.053 0.050 0.049 0.049
0.500 (38) 0.051 0.060 0.062 0.051 0.051 0.051
0.583 (71) 0.051 0.059 0.080 0.051 0.050 0.050
0.667 (131) 0.051 0.056 0.146 0.052 0.051 0.051
0.750 (241) 0.052 0.057 0.402 0.058 0.058 0.053
0.833 (443) 0.052 0.056 0.976 0.081 0.083 0.063
0.917 (815) 0.049 0.051 1.000 0.420 0.462 0.171
Figure S5: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n^{\tau}$ and $n=\sum_{i=1}^{3}500=1500$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S3.5 Robustness to misspecification for high dimensional setup
In general, all the approaches examined so far rely on the assumed normal model and are not guaranteed to be robust under model misspecification. We can assess numerically the robustness of the various competitors using simulations. We consider three different distributions for the true generating process: multivariate $t$, multivariate skew-normal (azzalini1999statistical) or multivariate Laplace. In more detail, a multivariate $t$ distribution with location $0_{p}$, scale matrix $I_{p}$ and degrees of freedom $5$, a multivariate skew-normal distribution with location $1_{p}$, scale matrix $\Omega=(\omega_{jl})=(0.2)^{|j-l|}$ and shape parameter $1_{p}$, and a multivariate Laplace distribution with mean vector $1_{p}$ and identity covariance matrix. Simulation results are based on $10,000$ replications.
For the homoscedastic case, Figures S6S7 show the empirical size at the nominal level $\alpha=0.05$ if the distribution is misspecified. We see that the directional test still maintains the highest accuracy.
Figure S6: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (11) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=100$.
Figure S7: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett corrected test (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2, respectively) for hypothesis (11) with $g=3$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$.
S3.6 Robustness to misspecification for moderate setup
In this section we investigate the robustness to misspecification. The true generating processes are multivariate $t$, multivariate skew-normal or multivariate Laplace distributions. Here we setup $p=n_{i}^{\tau}$ with $n_{i}\in\{100,500\}$.
In more detail, a multivariate $t$ distribution with location $0_{p}$, scale matrix $I_{p}$ and degrees of freedom $5$, a multivariate skew-normal distribution with location $1_{p}$, scale matrix $\Omega=(\omega_{jl})=(0.2)^{|j-l|}$ and shape parameter $1_{p}$, and a multivariate Laplace distribution with mean vector $1_{p}$ and identity covariance matrix.
For hypothesis (8) in the paper, Figures S8S9 and Tables S5S6 show the empirical size at the nominal level $\alpha=0.05$ if the underling distribution is misspecified. We see that the directional test still maintains the hightest accuracy.
Figure S8: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=100$ and $g=3$.
Figure S9: Empirical size for the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2, respectively) for hypothesis (8) in the paper, at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$ and $g=3$.
Table S5: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=3$, at nominal level $\alpha=0.05$.
True Distribution $\tau$ ($p$) DT CLT LRT BC Sko1 Sko2
Multivariate $t$ 0.250 (3) 0.050 0.069 0.053 0.050 0.049 0.049
0.333 (4) 0.044 0.063 0.048 0.044 0.042 0.042
0.417 (6) 0.048 0.065 0.054 0.048 0.047 0.047
0.500 (10) 0.047 0.059 0.055 0.046 0.045 0.045
0.583 (14) 0.050 0.060 0.064 0.050 0.049 0.048
0.667 (21) 0.049 0.061 0.075 0.049 0.048 0.048
0.750 (31) 0.044 0.054 0.096 0.045 0.043 0.043
0.833 (46) 0.042 0.049 0.148 0.043 0.041 0.040
0.917 (68) 0.042 0.047 0.302 0.047 0.045 0.041
Multivariate skew-normal 0.250 (3) 0.050 0.070 0.053 0.050 0.049 0.049
0.333 (4) 0.050 0.070 0.055 0.050 0.050 0.049
0.417 (6) 0.047 0.064 0.053 0.047 0.046 0.046
0.500 (10) 0.050 0.062 0.059 0.050 0.048 0.048
0.583 (14) 0.051 0.064 0.068 0.051 0.049 0.049
0.667 (21) 0.048 0.059 0.076 0.048 0.046 0.045
0.750 (31) 0.048 0.057 0.100 0.048 0.047 0.046
0.833 (46) 0.050 0.057 0.159 0.052 0.049 0.048
0.917 (68) 0.049 0.055 0.312 0.054 0.052 0.048
Multivariate Laplace 0.250 (3) 0.048 0.067 0.050 0.048 0.046 0.046
0.333 (4) 0.051 0.069 0.055 0.051 0.050 0.050
0.417 (6) 0.047 0.064 0.054 0.048 0.046 0.046
0.500 (10) 0.047 0.061 0.057 0.047 0.046 0.046
0.583 (14) 0.043 0.054 0.058 0.043 0.042 0.042
0.667 (21) 0.044 0.055 0.073 0.044 0.043 0.043
0.750 (31) 0.042 0.049 0.091 0.043 0.040 0.040
0.833 (46) 0.038 0.044 0.145 0.040 0.038 0.036
0.917 (68) 0.039 0.046 0.301 0.045 0.042 0.038
Table S6: Empirical size of the directional test (DT), central limit theorem test (CLT), likelihood ratio test (LRT), Bartlett correction (BC) and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (8) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=500$ and $g=3$, at nominal level $\alpha=0.05$.
True Distribution $\tau$ ($p$) DT CLT LRT BC Sko1 Sko2
Multivariate $t$ 0.250 (4) 0.052 0.071 0.052 0.052 0.051 0.051
0.333 (7) 0.055 0.069 0.057 0.055 0.055 0.055
0.417 (13) 0.050 0.062 0.052 0.050 0.050 0.050
0.500 (22) 0.052 0.063 0.057 0.052 0.051 0.051
0.583 (37) 0.048 0.055 0.058 0.048 0.048 0.048
0.667 (62) 0.044 0.051 0.068 0.044 0.044 0.044
0.750 (105) 0.045 0.049 0.099 0.045 0.045 0.044
0.833 (177) 0.043 0.047 0.212 0.045 0.045 0.043
0.917 (297) 0.043 0.046 0.618 0.052 0.052 0.046
Multivariate skew-normal 0.250 (4) 0.051 0.069 0.052 0.051 0.051 0.051
0.333 (7) 0.052 0.068 0.053 0.052 0.051 0.051
0.417 (13) 0.051 0.062 0.053 0.050 0.050 0.050
0.500 (22) 0.049 0.058 0.054 0.049 0.049 0.049
0.583 (37) 0.051 0.060 0.062 0.051 0.050 0.050
0.667 (62) 0.052 0.058 0.076 0.052 0.052 0.052
0.750 (105) 0.055 0.058 0.108 0.055 0.055 0.054
0.833 (177) 0.048 0.051 0.218 0.049 0.049 0.048
0.917 (297) 0.051 0.054 0.612 0.058 0.058 0.053
Multivariate Laplace 0.250 (4) 0.048 0.067 0.050 0.048 0.048 0.048
0.333 (7) 0.053 0.069 0.054 0.053 0.052 0.052
0.417 (13) 0.049 0.063 0.053 0.049 0.049 0.049
0.500 (22) 0.046 0.056 0.051 0.047 0.046 0.046
0.583 (37) 0.048 0.054 0.056 0.048 0.048 0.048
0.667 (62) 0.046 0.054 0.070 0.046 0.046 0.046
0.750 (105) 0.045 0.050 0.099 0.046 0.045 0.044
0.833 (177) 0.041 0.045 0.209 0.043 0.042 0.041
0.917 (297) 0.037 0.039 0.609 0.043 0.043 0.038
S4 Simulation studies for Heteroscedastic one-way MANOVA
This section is studied the performance of directional test for heteroscedastic one-way MANOVA, comparing with LRT, Sko1 and Sko2. In particular, when the number of groups $g=2$, we also consider the $F$-approximation for the Behrens-Fisher test $T^{*2}$. The simulation results are computed via Monte Carlo simulation based on 10,000 replications.
S4.1 Empirical results for the moderate setup
Groups of size $n_{i},i\in\{1,\dots,g\}$, are generated from a $p$-variate standard normal distribution $N_{p}(0_{p},\Lambda_{i}^{-1})$ under the null hypothesis. We use an autoregressive structure for the covariance matrices. i.e. $\Lambda_{i}^{-1}=(\sigma_{jl})_{p\times p}=(\rho_{i}^{|j-l|})_{p\times p}$, with the $\rho_{i}$ chosen to an equally-distance sequence from 0.1 to 0.9 of length $g$. For each simulation experiment, we show results for $p=\lceil n_{i}^{\tau}\rceil$ with $\tau=j/24$, $j\in\{6,7,\cdots,22\}$ and $n_{i}\in\{100,500,1000\}$ and $k=2$. Note that for $n_{i}=1000$ the simulations results are based on $5000$ replications when $j\in\{21,22\}$ due to the expensive computational cost.
Table S7: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100,500,1000$, at nominal level $\alpha=0.05$.
$n_{i}$ $\tau$ ($p$) DT BF LRT Sko1 Sko2
100 0.250 (4) 0.053 0.054 0.057 0.053 0.053
0.333 (5) 0.050 0.052 0.056 0.050 0.050
0.417 (7) 0.049 0.051 0.057 0.050 0.050
0.500 (10) 0.051 0.056 0.065 0.052 0.052
0.583 (15) 0.052 0.061 0.077 0.052 0.052
0.667 (22) 0.049 0.065 0.094 0.051 0.050
0.750 (32) 0.051 0.082 0.147 0.055 0.053
0.833 (47) 0.052 0.115 0.270 0.067 0.061
0.917 (69) 0.064 0.183 0.594 0.112 0.092
500 0.250 (5) 0.050 0.050 0.051 0.050 0.050
0.333 (8) 0.051 0.051 0.052 0.051 0.051
0.417 (14) 0.051 0.052 0.055 0.051 0.051
0.500 (23) 0.051 0.054 0.059 0.051 0.051
0.583 (38) 0.054 0.061 0.070 0.054 0.054
0.667 (63) 0.051 0.068 0.089 0.052 0.052
0.750 (106) 0.048 0.084 0.143 0.051 0.050
0.833 (178) 0.051 0.154 0.382 0.062 0.058
0.917 (298) 0.060 0.392 0.923 0.137 0.106
1000 0.250 (6) 0.052 0.052 0.053 0.052 0.052
0.333 (10) 0.050 0.050 0.051 0.050 0.050
0.417 (18) 0.046 0.048 0.048 0.046 0.046
0.500 (32) 0.052 0.055 0.058 0.052 0.052
0.583 (57) 0.052 0.058 0.063 0.052 0.052
0.667 (100) 0.050 0.066 0.083 0.050 0.050
0.750 (178) 0.047 0.088 0.154 0.050 0.049
0.833 (317) 0.051 0.179 0.459 0.065 0.059
0.917 (563) 0.064 0.528 0.987 0.155 0.112
Figure S10: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (12) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to $n_{i}=1000$, respectively.
S4.2 Empirical results for large number $g$ of groups
Table S8: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$.
$\tau$ ($p$) DT LRT Sko1 Sko2
0.250 (4) 0.050 0.081 0.050 0.050
0.333 (5) 0.050 0.091 0.050 0.050
0.417 (7) 0.049 0.122 0.050 0.049
0.500 (10) 0.050 0.186 0.051 0.050
0.583 (15) 0.054 0.391 0.058 0.054
0.667 (22) 0.047 0.751 0.056 0.048
0.750 (32) 0.052 0.990 0.087 0.062
0.833 (47) 0.049 1.000 0.223 0.095
0.917 (69) 0.048 1.000 0.932 0.405
Figure S11: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=30$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
Table S9: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=1000$ and $g=5$, at nominal level $\alpha=0.05$
$\tau$ ($p$) DT LRT Sko1 Sko2
0.250 (6) 0.053 0.055 0.053 0.053
0.333 (10) 0.055 0.058 0.055 0.055
0.417 (18) 0.048 0.054 0.048 0.048
0.500 (32) 0.048 0.061 0.048 0.048
0.583 (57) 0.051 0.089 0.052 0.052
0.667 (100) 0.054 0.160 0.055 0.054
0.750 (178) 0.053 0.462 0.058 0.055
0.833 (317) 0.050 0.983 0.083 0.063
Figure S12: Empirical size of the directional test (DT), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper, with $p=n_{i}^{\tau}$ and $n_{i}=1000$ and $g=5$, at nominal level $\alpha=0.05$ given by the gray horizontal line.
S4.3 Robustness to misspecification
Table S10: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=100$ and $g=2$, at nominal level $\alpha=0.05$.
Distribution $\tau$ ($p$) DT BF LRT Sko1 Sko2
Multivariate $t$ 0.250 (4) 0.048 0.048 0.052 0.048 0.048
0.333 (5) 0.047 0.047 0.053 0.047 0.047
0.417 (7) 0.048 0.048 0.055 0.048 0.048
0.500 (10) 0.050 0.049 0.060 0.050 0.050
0.583 (15) 0.045 0.044 0.064 0.045 0.045
0.667 (22) 0.047 0.046 0.078 0.047 0.047
0.750 (32) 0.042 0.040 0.102 0.043 0.042
0.833 (47) 0.041 0.038 0.166 0.044 0.042
0.917 (69) 0.043 0.038 0.375 0.054 0.048
Multivariate skew-normal 0.250 (4) 0.046 0.046 0.051 0.046 0.046
0.333 (5) 0.052 0.052 0.058 0.052 0.052
0.417 (7) 0.053 0.053 0.061 0.053 0.053
0.500 (10) 0.054 0.054 0.064 0.054 0.054
0.583 (15) 0.049 0.049 0.070 0.050 0.049
0.667 (22) 0.048 0.047 0.080 0.048 0.048
0.750 (32) 0.050 0.048 0.110 0.051 0.050
0.833 (47) 0.052 0.051 0.187 0.056 0.053
0.917 (69) 0.048 0.048 0.379 0.061 0.053
Multivariate Laplace 0.250 (4) 0.049 0.049 0.054 0.049 0.049
0.333 (5) 0.044 0.044 0.050 0.044 0.044
0.417 (7) 0.045 0.045 0.051 0.045 0.045
0.500 (10) 0.046 0.046 0.055 0.046 0.045
0.583 (15) 0.045 0.045 0.063 0.045 0.045
0.667 (22) 0.044 0.043 0.075 0.044 0.044
0.750 (32) 0.040 0.039 0.097 0.042 0.040
0.833 (47) 0.039 0.038 0.169 0.043 0.040
0.917 (69) 0.034 0.032 0.372 0.044 0.038
Table S11: Empirical size of the directional test (DT), Behrens-Fisher test (BF) (nel1986), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) in the paper with $p=n_{i}^{\tau}$ and $n_{i}=500$ and $g=2$, at nominal level $\alpha=0.05$.
Distribution $\tau$ ($p$) DT BF LRT Sko1 Sko2
Multivariate $t$ 0.250 (5) 0.050 0.050 0.052 0.050 0.050
0.333 (8) 0.049 0.049 0.050 0.049 0.049
0.417 (14) 0.047 0.047 0.050 0.047 0.047
0.500 (23) 0.045 0.045 0.050 0.045 0.045
0.583 (38) 0.051 0.051 0.063 0.051 0.051
0.667 (63) 0.044 0.044 0.070 0.044 0.044
0.750 (106) 0.044 0.043 0.108 0.044 0.044
0.833 (178) 0.044 0.043 0.236 0.048 0.046
0.917 (298) 0.047 0.045 0.705 0.064 0.053
Multivariate skew-normal 0.250 (5) 0.051 0.051 0.052 0.051 0.051
0.333 (8) 0.051 0.051 0.052 0.051 0.051
0.417 (14) 0.055 0.055 0.059 0.055 0.055
0.500 (23) 0.054 0.054 0.060 0.054 0.054
0.583 (38) 0.051 0.051 0.061 0.051 0.051
0.667 (63) 0.049 0.049 0.074 0.049 0.049
0.750 (106) 0.051 0.051 0.117 0.052 0.052
0.833 (178) 0.048 0.048 0.242 0.054 0.050
0.917 (298) 0.052 0.051 0.699 0.070 0.058
Multivariate Laplace 0.250 (5) 0.054 0.054 0.054 0.053 0.053
0.333 (8) 0.053 0.053 0.054 0.053 0.053
0.417 (14) 0.050 0.050 0.052 0.050 0.050
0.500 (23) 0.051 0.051 0.056 0.051 0.051
0.583 (38) 0.047 0.047 0.056 0.047 0.047
0.667 (63) 0.046 0.046 0.070 0.047 0.046
0.750 (106) 0.042 0.042 0.106 0.043 0.043
0.833 (178) 0.040 0.040 0.235 0.043 0.040
0.917 (298) 0.033 0.032 0.708 0.051 0.040
Figure S13: Empirical size for the directional test (DT), Behrens-Fisher test (BF), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9) with $g=2$, at nominal level $\alpha=0.05$ given by the dashed gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions if the true generating process, respectively, with $n_{i}=100$.
Figure S14: Empirical size for the directional test (DT), Behrens-Fisher test (BF), likelihood ratio test (LRT), and two skovgaard:2001’s modifications (Sko1 and Sko2) for hypothesis (9), at nominal level $\alpha=0.05$ given by the gray horizontal line. The left, middle and right panels correspond to multivariate $t$, multivariate skew-normal, and multivariate Laplace distributions of the true generating process, respectively, with $n_{i}=500$ and $g=2$.
S5 Application: Microarray Data of Breast Cancer Patients
The dataset studied is taken from transcript profiles of primary breast tumors to investigate the p53 status. It is based on 251 samples predominantly containing tumor tissue that underwent sequencing to identify the p53 mutation status through RNA extraction and subsequent microarray analysis. The RNA was extracted from each sample by using Affymetrix U133 oligonucleotide microarrays, targeting over 30,000 genes (miller2005expression). Here we focus on a subset of genes corresponding to a signature that is considered important in breast cancer from the research of bonavita2020antagonistic, which stores the appropriately normalized expression values of $p=22$ gene products. The samples were divided into three groups based on tumor grade. After excluding two samples due to the missing grade variable, we obtain $g=3$ groups with $n_{1}=67$, $n_{2}=128$ and $n_{3}=54$ samples for grade I, grade II and grade III, respectively. Within a high dimensional framework, we have $p/\min(n_{i})=0.407$. Homoscedastic one-way MANOVA leads to the following $p$-values: $1.565\times 10^{-7}$ for the directional test, $1.966\times 10^{-12}$ for the central limit theorem test, $2.627\times 10^{-8}$ for the likelihood ratio test, $1.915\times 10^{-7}$ for the Bartlett corrected test, and $1.828\times 10^{-7}$ and $1.877\times 10^{-7}$ for the two Skovgaard’s modifications. All methods clearly indicate to reject the null hypothesis of equal breast tumors gene expression signature in case of homoscedastic groups. As already seen, the likelihood ratio test results in the smallest $p$-value, whereas its Skovgaard’s modified versions and the Bartlett corrected test are closer to the exact directional $p$-value, consistent with our simulations. To check the reliability of the various methods in the homoscedastic case, another simulation study on the gene expression data can be conducted. The empirical size based on $10,000$ replications for the directional, central limit theorem test, likelihood ratio, Bartlett corrected and two Skovgaard’s modifications tests are 0.049, 0.060, 0.087, 0.049 0.048 and 0.047, respectively. Those of central limit theorem test and likelihood ratio test are a little inflated. The remaining approaches, including the directional test, appear to be accurate.
Assuming heteroscedastic groups, for hypothesis (9) the directional $p$-value is $2.861\times 10^{-5}$, the $p$-value of the likelihood ratio test is $1.928\times 10^{-7}$, and the $p$-values of the two Skovgaard’s modifications are $1.772\times 10^{-5}$ and $2.079\times 10^{-5}$. The same conclusion is hence drawn by all methods, but once again the $p$-value obtained from the likelihood ratio test is relatively smaller than the others. Such discrepancies are particularly relevant in biostatistical problems, and more generally in medical research, where it is important to look at the specific size of $p$-values to get an idea about the clinical implications and relevance of the findings for the population of patients of interest. Even under heterescedasticity a simulation study can be performed. The empirical sizes of the directional test, the likelihood ratio test and the two Skovgaard’s modifications based on $10,000$ replications are 0.051, 0.181, 0.057 and 0.054, respectively. The directional test performs very well, as expected, while the likelihood ratio test seems still not reliable.